### v1

# PatentScore: Multi-dimensional Evaluation of LLM-Generated Patent Claims

Abstract
Natural language generation (NLG) metrics play a central role in evaluating generated texts, but are not well suited for the structural and legal characteristics of patent documents. Large language models (LLMs) offer strong potential in automating patent generation, yet research on evaluating LLM-generated patents remains limited, especially in evaluating the generation quality of patent claims, which are central to defining the scope of protection. Effective claim evaluation requires addressing legal validity, technical accuracy, and structural compliance. To address this gap, we introduce PatentScore, a multi-dimensional evaluation framework for assessing LLM-generated patent claims. PatentScore incorporates: (1) hierarchical decomposition for claim analysis; (2) domain-specific validation patterns based on legal and technical standards; and (3) scoring across structural, semantic, and legal dimensions. Unlike general-purpose NLG metrics, PatentScore reflects patent-specific constraints and document structures, enabling evaluation beyond surface similarity. We evaluate 400 GPT-4o-mini generated Claim 1s and report a Pearson correlation of $r=0.819$ with expert annotations, outperforming existing NLG metrics. Furthermore, we conduct additional evaluations using open models such as Claude-3.5-Haiku and Gemini-1.5-flash, all of which show strong correlations with expert judgments, confirming the robustness and generalizability of our framework.
1 Introduction
Patents involve rapidly increasing, significant technological, economic, and social impacts. More than 3 million patent applications were filed globally in 2021 (WIPO, 2022), representing millions of dollars in the intellectual property (IP) market (OECD, 2021). However, pattern drafting, reviewing, and filing are not only time and financially costly but also technically challenging due to their complex structures, legal constraints, and technical specifications (Surden, 2019). Drafting a patent application could cost 30-40 hours of work by patent attorneys and technical experts, costing USD\$5,000 to \$15,000 (WIPO, 2022). Therefore, automating patent generation and examination has a growing global urgency and priority (Ciccatelli, 2019), where AI, specifically natural language generation (NLG) and large language models (LLMs), play essential roles (Aristodemou and Tietze, 2018). This demand grows rapidly over time, as illustrated in Figure 1 of patent applications and registrations in Australia from 2011 to 2021. Widely required in the global patent domain, further emphasizing the importance of patent automation (Fujii, 2007).
Figure 1: Patent applications and registrations in Australia in 2011-2021 (IP Australia, 2022).
This trend of increasing patent activities is even more pronounced in major economies. China and the United States lead global patent applications (WIPO, 2022). The high application volume at these offices has led to increasing backlogs and processing times, emphasizing the need for AI-assisted patent automation (Lee and Hsiang, 2020). Patent automation has the potential to revolutionize the patent filing process, significantly reducing the time, cost, and human effort involved, and significantly enhancing the efficiency of patent filing and innovation processes. (Ebrahim, 2016).
A specific area concerned in this paper is the automated evaluation of LLM-generated patent Claim 1, which is the most important section in patent applications. Claim 1 defines comprehensive but precise legal and technical scopes of IP value and protection and determines critical legal boundaries of patent rights (Merges, 1997). It is essential to ensure legal and technical accuracy when evaluating AI-generated patent claims. Errors or ambiguities in Claim 1 can lead to litigation, invalidation, or reduced protection, making precise evaluation of generated patent claims critically important (Bessen and Meurer, 2008). Addressing these unique requirements and challenges demands specialized evaluation frameworks for LLM-generated patent claims.
However, this demand cannot be met by existing NLG-based evaluation metrics originally designed for general text evaluation, which focus primarily on lexical and semantic similarities (Papineni et al., 2002; Lin, 2004). For example, BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004) primarily evaluate n-gram overlap, and BERTScore (Zhang et al., 2019) focuses on contextual embeddings for semantic similarity. They inherently lack the mechanisms to address critical patent-specific requirements such as legal compliance and technical consistency, and fail to address patent-specific requirements such as structural coherence, claim dependency, antecedent basis, and legal compliance, which are crucial for evaluating patent claims (Merges et al., 1997). In addition, while LLMs have demonstrated substantial progress in transforming many language tasks such as question-answering and machine vision Brown et al. (2020), their applications in generating and evaluating patent claims are unexplored.
Accordingly, we address these limitations by integrating claim structural validation and claim-specific legal criteria into a multi-dimensional claim evaluation framework PatentScore, significantly enhancing claim evaluation accuracy and reliability. Compared with NLG metrics shown in Figure 2 (a), which only analyzes text-level similarity without considering their hierarchical structures or legal specifications, PatentScore shown in Figure 2 (b) systematically analyzes hierarchical claim structures by decomposing claim structures and integrates patent-specific legal validation patterns into effectively evaluating aspects like claim dependency and ambiguous scope. These ensure legal and technical accuracy of generated claims, thus significantly enhancing the reliability of patent claim evaluation.
Figure 2: The comparison between NLG metric and PatentScore.
Illustrated on GPT-4o-mini-generated patent claims, PatentScore provides an effective pattern generation evaluation framework that extends the Chain-of-Thought (Wei et al., 2022) to structural, legal, and semantic aspects of patent claims. It develops and integrates the following three important aspects into LLM-generated claim evaluation:
  • Patent-structured analysis architecture: A novel framework that systematically analyzes patent claims through hierarchical decomposition, validating structural and technical compliance. It incorporates patent-specific requirements such as claim structure and scope verification, essential for legal validity yet overlooked by existing NLG metrics.
  • Patent-specific quality metrics: A comprehensive evaluation introducing new metrics for legal and technical accuracy, leveraging semantic and terminological consistency in patents.
  • Verifiable evaluation protocol: A systematic validation demonstrating a strong correlation with expert judgments (r = 0.819) through a rigorous evaluation at the component level of seven key patent-oriented factors.
Unlike prior studies that primarily focused on novelty assessment or prior art retrieval, this paper addresses a distinct yet critical problem: evaluating the intrinsic quality of individual claims generated by LLMs. As the legal foundation of patent protection, claims demand structural, technical, and legal precision. To this end, we propose PatentScore, a multi-dimensional evaluation framework, and conduct large-scale experiments on claims generated by GPT-4o-mini. The results show a strong correlation with expert judgments ($r=0.819$) and demonstrate that PatentScore significantly outperforms existing metrics such as BLEU, ROUGE, GPTScore, and cosine similarity. PatentScore provides a reliable and generalizable framework for evaluating AI-generated patent claims, serving as a practical foundation for building LLM-based pipelines for automated patent drafting and assessment.
2 Background & Related Work
2.1 Specifications in Patent Claims
Claim 1 (Example)
1. An extendible trailer tongue comprising:
    a first tubular member;
    a second tubular member telescopically received within the first tubular member; and
    a locking mechanism to secure the second tubular member in a selected position relative to the first tubular member.
Table 1: An example of Claim 1 from the patent titled "Multiple component headgear system" (Siprut, 2000).
Patents are specialized legal-technical documents that protect IP rights. Among the various components of a patent, claims define the legal boundaries of IP protection, with Claim 1 serving as the cornerstone. Claim 1 establishes the broadest scope of protection and directly impacts the patent’s validity and enforceability (WIPO, 2023; Merges et al., 1997).
Unlike general technical documents, patent claims require adherence to strict structural formats, standardized terminology, and legal rigor. Table 1 illustrates a well-formed Claim 1, which demonstrates these key requirements:
  • Structural format: Compliance with strict formatting requirements.
  • Technical clarity: Precise definition of technical elements and relationships.
  • Terminology consistency: Use of standardized patent-specific language.
2.2 Patent Generation and Evaluation
Automating patent generation and evaluation has become an emerging agenda, yet remains underexplored. Large language models (LLMs) have recently been applied to generate technical and legal content Brown et al. (2020), showing promise in zero-shot reasoning Kojima et al. (2023) and complex instruction following Li et al. (2024). However, most existing studies focus on general text generation, without addressing the strict structural and legal requirements of patent claims Surden (2019). Patent evaluation involves both linguistic and domain-specific dimensions. Traditionally, especially Claim 1,are reviewed manually by examiners and attorneys, which is time-consuming and inconsistent. As patent applications increase globally, automated evaluation systems must assess technical correctness, legal soundness, and structural compliance simultaneously (Surden, 2019).
Yet, evaluating LLM-generated patent claims is particularly challenging due to their multi-faceted nature. Minor errors in structure, terminology, or legal scope can result in litigation or invalidation. Existing evaluation metrics such as BLEU, ROUGE, and BERTScore Sellam et al. (2020), primarily assess linguistic similarity but fail to capture critical patent-specific elements like claim scope, legal consistency, and structural validity.
Recent LLM evaluation frameworks, including PromptScore, HolisticEval, and LMdiff (Gao et al., 2024; Wang et al., 2023; Liu et al., 2024), provide task-specific coherence assessments. However, they do not account for the legal and structural constraints inherent to patents. PatentScore addresses this gap by incorporating patent-specific evaluation dimensions into its framework.
2.3 NLG Metrics for Patent Evaluation
NLG metrics dominate patent generation evaluation, we review them specifically. First, traditional n-gram based metrics like BLEU and ROUGE measure lexical overlap between generated and reference texts. However, these lexical similarities overlook rich semantic relationships and domain specialties Sellam et al. (2020), which are crucial particularly for specialized contexts like patent claims with legal and technical precision Surden (2019). Then, metrics based on deep models like Transformer enable more nuanced evaluation. For example, BERTScore utilizes contextual embeddings to assess semantic similarity, while BLEURT Sellam et al. (2020) combines BERT’s representational power with task-specific optimization. GPTScore Fu et al. (2023) further leverages LLMs to evaluate semantic coherence, logical consistency, and contextual appropriateness, making it promising for legal and technical content evaluation.
Despite these advancements, the NLG/LLM evaluation metrics struggle to address unique challenges posed by patent documents with strong legal features and requirements. Unlike general text, patent claims require rigorous structural compliance, precise legal terminology, and domain-specific validation. Existing metrics fall short in assessing critical elements such as claim scope definition, antecedent basis establishment, and hierarchical relationships between claims. This highlights the necessity for specialized evaluation frameworks capable of addressing both linguistic and patent-specific requirements comprehensively.
3 PatentScore for Patent Claim Evaluation
Figure 3: The PatentScore Framework.
3.1 The PatentScore Framework for multi-dimensional patent claim evaluation.
PatentScore is a systematic evaluation framework for patent claims that integrates structural, legal, and semantic analyses in a decomposed manner. Figure 3 illustrates how these components work together to generate overall quality scores while maintaining independence in their evaluations.
First, the structural analysis identifies three key characteristics: claim structure ($M_{CS}$), punctuation pattern ($M_{CP}$), and antecedent basis ($M_{AB}$). Each is scored on a scale of 1-5. Then, the legal analysis examines element referencing ($M_{ER}$), validity/uniqueness ($M_{VU}$), and ambiguous scope ($M_{AS}$), also scored on a scale of 1-5. For semantic analysis, we employ BERTScore ($M_{BS}$), a proven metric in NLG, which provides scores between 0 and 1, later scaled to 1–5 for consistency with other metrics. To systematically evaluate these components, we leverage GPT-4o-mini with carefully designed structural, legal and semantic prompts for them. Each component is assessed through specialized prompts that guide the model to focus on specific aspects below. Detailed prompts are provided in Appendix A.
Structural component
evaluates the technical composition and formatting of claims, following standard patent writing conventions and WIPO patent drafting guidelines WIPO (2023).
Legal component
analyzes compliance with patent law requirements and clarity standards based on USPTO’s MPEP guidelines USPTO (2023) and WIPO patent drafting guidelines WIPO (2023).
Semantic component
employs BERTScore to measure semantic similarity, a proven approach in NLG for assessing textual coherence.
These components apply specific prompts that enable PatentScore to perform targeted evaluation of each metric while maintaining consistency across assessments. PatentScore enables decomposing the complex task of patent claim evaluation into well-defined sub-tasks that can be systematically analyzed using LLMs. To mitigate the inherent variability of GPT model outputs, we perform ten iterations of evaluation for each metric and use the mean score as the final value. To further enhance output consistency, we fix the temperature parameter at 0.3. This configuration ensures more reliable and stable evaluation results. Furthermore, we validate the robustness and generalizability of PatentScore by applying the same evaluation pipeline to multiple LLMs, including GPT-4o-mini, Claude-3.5-Haiku, and Gemini-1.5-flash. Since these evaluation criteria are grounded in internationally recognized guidelines (e.g., WIPO (WIPO, 2023), USPTO (USPTO, 2023)), PatentScore can be readily adapted to other jurisdictions with minimal adjustments.
3.2 Decomposed Patent Evaluation Components
We propose patent evaluation components in structural, legal, and semantic aspects, respectively.
Structural Components consist of claim structure, claim punctuation, and antecedent basis.
Claim Structure ($M_{CS}$): Assesses the preamble, transitional phrases, and elements to ensure structural completeness and clear enforceability.
Claim Punctuation ($M_{CP}$): Focuses on critical punctuation marks (e.g., colons, semicolons, "and," periods) to maintain compliance with patent drafting standards and ensure accurate interpretation.
Antecedent Basis ($M_{AB}$): Evaluates the proper establishment of antecedent basis by checking article consistency in accordance with patent drafting standards.
Legal Components comprise element referencing, validity and uniqueness, and ambiguous scope.
Element Referencing ($M_{ER}$): Verifies the accuracy and consistency of references to technical components, ensuring clear relationships and dependencies among claim elements.
Validity and Uniqueness ($M_{VU}$): Evaluates whether claim elements exhibit technically distinctive functions or expressive originality beyond common domain knowledge, based on prompt-guided assessment.
Ambiguous Scope ($M_{AS}$): Identifies and resolves overly broad or ambiguous language, ensuring the claim scope is precise and legally enforceable.
Semantic Components consist of BERTScore.
BERTScore ($M_{BS}$): Measures semantic similarity and contextual alignment of claims with reference descriptions, ensuring accurate representation of domain-specific terminology.
3.3 Prompt Design for Claim Evaluation
In this study, we propose a LLM-adaptable prompt framework for systematic evaluation of patent claims. Our prompt engineering methodology builds upon Chain of Thought(CoT) (Wei et al., 2022) reasoning, extending its sequential inference process to accommodate the multidimensional characteristics of patent evaluation. Rather than implementing a simple step-by-step analysis, we develop an integrated evaluation framework that considers both the interdependencies among structural elements within patent claims and the hierarchical relationships of legal requirements.
This expanded prompt structure ensures consistency across evaluations while enabling parallel processing of multiple evaluation dimensions, allowing for a more sophisticated and systematic analysis. Specifically, building upon the structured reasoning principles of CoT, the prompts are tailored to extend claim analysis and scoring processes to meet the structural, legal, and semantic demands of patents. At each stage, explicit reasoning steps and evaluation criteria are incorporated, significantly enhancing the reliability and reproducibility of the evaluation process.
Component-specific Template Design We designed a standardized base template inspired by CoT, extending its sequential reasoning to a multidimensional framework tailored for patent claim evaluation, enabling both consistency and comprehensive analysis.
TASK: Patent Claim 1 Assessment
FOCUS: Component-specific evaluation
OUTPUT FORMAT: Detailed analysis and score (1-5)
I. Evaluation Criteria
Primary Requirements:
- [Component-specific requirement 1]
- [Component-specific requirement 2]
Secondary Requirements:
- [Additional evaluation points]
II. Analysis Procedure
1. Identify relevant claim elements
2. Evaluate against requirements
3. Assign score with justification
III. Scoring Rubric
5: Exceeds all requirements with exceptional clarity
4: Meets all requirements effectively
3: Meets primary requirements adequately
2: Partially meets requirements with deficiencies
1: Fails to meet critical requirements
Implementation by Components We customize the templates for structural analysis ($M_{CS}$, $M_{CP}$, $M_{AB}$) and legal compliance ($M_{ER}$, $M_{VU}$, $M_{AS}$) components. For semantic evaluation ($M_{BS}$), we utilize BERTScore, which provides an automated assessment of semantic similarity between generated and reference claims through contextual embeddings. This structured approach, combining specialized prompts for structural and legal assessments with automated semantic evaluation, enables a reliable and reproducible evaluation of patent claim quality across all dimensions of our framework. Detailed prompts for each evaluation component are provided in Appendix A
3.4 Evaluation Process
PatentScore evaluates patent claims through four sequential stages inspired by but expanding the CoT approach, enabling systematic quality evaluation. The evaluation process consists of four core stages: (1) Component decomposition breaks down patent claims into analyzable elements; (2) Component analysis includes content analysis (identifying and mapping claim elements), structure validation (verifying structural correctness), standard comparison (evaluating against patent requirements) and score calculation (calculating component-specific scores); (3) Semantic analysis applies BERTScore ($M_{BE}$); and (4) Score integration combines scores using ablation-derived weights.
The weights are determined by two complementary measures: correlation coefficient ($r$) with expert assessment, and mean absolute error (MAE) with expert scoring. To calculate performance changes, we conduct ablation experiments by removing each component individually and measuring the resulting changes in correlation ($\Delta r_{i}$) and MAE ($\Delta MAE_{i}$) compared to the full model. These measures reflect the contribution and degradation of each component to the overall model performance. The final weight for each component is computed as:
$w_{i}=\frac{\Delta r_{i}+\Delta MAE_{i}}{\sum_{j=1}^{7}(\Delta r_{j}+\Delta MAE_{j})}$ (1)
where $\Delta r_{i}=r_{full}-r_{i}$ and $\Delta MAE_{i}=MAE_{i}-MAE_{full}$ represent performance degradation when the $i$-th component is removed. To ensure comparability, all metrics ($Mi$) are normalized to a consistent range of [1, 5] before aggregation. The final PatentScore is then calculated as:
$\textit{PatentScore}=\sum_{i=1}^{7}w_{i}\cdot M_{i},$ (2)
where $M_{i}$ represents each evaluation metric, and $w_{i}$ is the corresponding weight.
Our PatentScore framework provides a systematic approach with a carefully structured process for patent claim evaluation. The evaluation algorithm decomposes patent claims and analyzes them through multiple perspectives including structural correctness, legal compliance, and semantic coherence. The framework’s distinctive feature lies in its weight determination mechanism, which optimizes the contribution of each component through comprehensive ablation studies. This data-driven approach to weight optimization ensures that the final evaluation score reflects the relative importance of each metric based on empirical evidence from expert evaluation. By combining structured analysis with semantic assessment and incorporating expert knowledge through weights, PatentScore offers a robust and comprehensive method for evaluating patent claim quality.
4 Data and Verification Measures
4.1 Claim Data Generation and Composition
We use a subset of the HUPD dataset (Suzgun et al., 2022), consisting of patents filed in 2016 and 2017. This subset includes the first claim of patents classified under Section A (Human Necessities) and Section G (Physics) of the International Patent Classification (IPC), comprising 400 pairs (200 from each section). The dataset was curated to support the comparative evaluation of patent claims generated by GPT-4o-mini and those written by human experts. Table 2 presents an example entry, which serves as the basis for assessing the legal and technical completeness of generated claims, and for validating our proposed sequential evaluation metric. Each entry includes the patent ID, the original human-authored Claim 1, and the corresponding GPT-4o-mini-generated version. The dataset is publicly available.(Our dataset will be available upon acceptance).
Patent Number: US20170123456
Gold-Claim A method for processing biometric data, comprising: receiving sensor data; extracting features; and authenticating a user.
LLM-Claim A method for biometric processing, comprising: obtaining sensor data; processing features; and performing authentication.
Expert Scores E1: 3.0  E2: 4.0  E3: 3.0 Mean Score: 3.33
Table 2: An example of patent claim data generated by LLMs and humans, with expert annotations.
4.2 Expert Evaluation
To verify the reliability of the LLM-generated claims in the dataset and the validity of the evaluation results, a panel of three experts was formed, consisting of one legal expert and two technical experts (PhD candidates). Each expert independently assessed the legal and technical completeness of all claims, and the evaluation was conducted using a Likert scale ranging from 1 to 5. As illustrated in Table 2, we also provide the scores made by three experts, i.e., E1-E3, and their mean score.
4.3 Verification Measures
To validate the effectiveness of PatentScore, we conducted a comparative analysis with standard NLG metrics, including BLEU, ROUGE-L, GPTScore, and Cosine Similarity. Evaluations across structural, legal, and semantic dimensions showed that PatentScore yields more consistent and fine-grained results. We further assessed the reliability of LLM-generated claims by comparing them with expert annotations using statistical measures such as Pearson correlation, Cronbach’s $\alpha$, and the Intraclass Correlation Coefficient (ICC), confirming the objectivity and consistency of the evaluation framework.
5 Evaluation Results
5.1 Claim Generation Quality Analysis
Metric $r$ $\rho$ $\tau$ MAE
PatentScore .819 .813 .665 .568
PatentScore Avg. .812 .793 .551 .512
BERTScore -.161 -.163 -.130 1.975
GPT Score .013 -.004 -.003 1.408
BLEU -.117 -.089 -.065 1.744
ROUGE-L -.159 -.112 -.080 2.499
Cosine Sim -.050 .030 .024 1.946
Table 3: The comparison of various metrics using Pearson correlation $r$, Spearman correlation $\rho$, Kendall’s Tau $\tau$ and MAE.
Our framework demonstrates strong alignment with human expert evaluation, achieving a Pearson correlation coefficient of $r=0.819$ ($p<0.01$) across all metrics, substantially surpassing existing methods. It is noteworthy that the consistent negative correlation shown by traditional metrics such as BERTScore ($r=-0.161$), ROUGE-L ($r=-0.159$), and BLEU ($r=-0.117$), indicating their limitations in patent claim evaluation. The effectiveness of the framework is further validated by its strong performance across multiple correlation measures, including Spearman’s rank correlation ($\rho=0.813$) and Kendall’s rank correlation ($\tau=0.665$). Beyond correlation coefficients, our approach demonstrates significantly lower error rates ($MAE=0.568$) compared to existing metrics, which exhibit substantially higher error values (e.g., MAE=2.499 for ROUGE-L).
Our results reveal that commonly used metrics show limited effectiveness in capturing the nuances of patent claim quality, as indicated by their negative correlations with expert judgments. While metrics like BERTScore and ROUGE-L show negative or near-zero correlation with expert scores, PatentScore exhibits strong positive correlation, indicating its suitability for the legal and structural nuances of patent claims. The consistently superior performance of our framework across diverse criteria establishes it as a robust and reliable method for patent claim assessment. To assess generalizability, we also evaluate PatentScore using open-source models, including Claude-3.5-Haiku and Gemini-1.5-Flash, which achieved Pearson correlations of 0.745 and 0.731 respectively demonstrating high alignment with expert annotations across models.
5.2 Quality Comparison of Expert Evaluation on Generated Claims
The Pearson correlation coefficients between the ratings by three experts range from 0.795 to 0.847, demonstrating their high correlation. Cronbach’s $\alpha$ is 0.931, indicating a very high level of internal consistency in expert evaluations. Furthermore, the ICC(3,k) value of 0.928 confirms excellent agreement among the evaluators, while Krippendorff’s $\alpha$ is 0.784, reflecting a substantial level of agreement. These findings statistically validate that the proposed evaluation framework captures rigorous and consistent criteria, ensuring high levels of reliability and reproducibility.
5.3 Ablation Study
Model Corr. ($r$) MAE Perf. Drop (%) Weight
Baseline ($\ddagger$) 0.818 0.568 - 1.0
- $M_{CS}$ 0.735 0.675 10.1 (18.8) 0.166
- $M_{CP}$ 0.723 0.667 11.6 (17.4) 0.171
- $M_{AB}$ 0.731 0.671 10.6 (18.1) 0.167
- $M_{ER}$ 0.734 0.670 10.3 (18.0) 0.163
- $M_{VU}$ 0.742 0.674 9.3 (18.7) 0.159
- $M_{AS}$ 0.729 0.677 10.9 (19.2) 0.173
- $M_{BS}$ 0.817 0.569 0.1 (0.2) 0.002
Notes: Numbers in parentheses indicate MAE increases. The baseline model ($\ddagger$) uses equal weights for all components.
Table 4: The comparison of metric contributions via ablation analysis.
We conducted an ablation study to analyze the impact of each metric on the overall framework performance, as shown in Table 4. The baseline model, which assigns equal weights (i.e., $w_{i}=1/7$) to all seven components, achieves a correlation of 0.818 with MAE of 0.568. Further, component-wise analysis reveals:
  • Structural metrics: Removing claim structure $M_{CS}$, punctuation pattern $M_{CP}$, and antecedent basis $M_{AB}$ leads to correlation drops of 10.1%, 11.6%, and 10.6%, respectively.
  • Legal metrics: Excluding element referencing $M_{ER}$, validity/uniqueness $M_{VU}$, and ambiguous scope $M_{AS}$ causes decreases of 10.3%, 9.3%, and 10.9%, respectively.
  • Semantic metric: Removing BERTScore $M_{BS}$ shows minimal impact of 0.1% decrease, highlighting its complementary role.
Following the final weights in Eqn. 1, we can calculate the patent score per Eqn. 2, where each weight $W$ reflects the relative importance of corresponding metrics derived from the ablation study.
PatentScore= $\begin{split}\textit{PatentScore}=&\,0.166\cdot M_{CS}+0.171\cdot M_{CP}+\\ &\,0.163\cdot M_{ER}+0.167\cdot M_{AB}+\\ &\,0.159\cdot M_{VU}+0.173\cdot M_{AS}+\\ &\,0.002\cdot M_{BS}\end{split}$ (3)
These findings highlight the minimal impact of BERTScore on overall framework performance. Removing BERTScore results in only a 0.1% decrease, indicating that semantic metrics play a complementary rather than critical role in patent claim evaluation. In contrast, structural and legal metrics demonstrate substantial influence, with performance drops of 9.3% to 11.6% when excluded underscoring the importance of structural validity and legal compliance in determining claim quality. While modern LLMs exhibit strong semantic consistency and accuracy, they still require explicit guidance to meet structural and legal requirements. This balance of semantic strength and structural/legal support affirms the necessity of a comprehensive evaluation framework like PatentScore.
5.4 Key Findings
Limitations of Existing Metrics: Conventional text evaluation metrics (BLEU, ROUGE-L, and BERTScore) demonstrate negative correlations ($r=-0.117\sim-0.161$) in patent claim evaluation. BLEU’s n-gram focus neglects logical dependencies and ROUGE-L prioritizes surface-level similarity over domain-specific coherence.
Balanced Component Evaluation: Our ablation study reveals comparable weights for structural ($0.166-0.171$) and legal components ($0.159-0.173$). The 10.6% performance drop from omitting antecedent basis metric ($M_{AB}$) exemplifies the necessity of multi-faceted evaluation approaches.
Semantic Analysis Role: While BERTScore shows limited independent impact (weight: $0.002$), it enhances the framework by validating domain-specific terminology consistency, particularly for emerging technology terms.
6 Conclusion
This study presents PatentScore, a multi-dimensional framework for evaluating LLM-generated patent claims. By integrating structural, legal, and semantic metrics tailored to patent language, it closely aligns with expert judgments (Pearson $r=0.819$) and outperforms standard NLG metrics. Patent claims require a blend of technical, legal, and linguistic elements—often missed by general-purpose metrics. PatentScore addresses this through a modular design that aligns language modeling with legal evaluation, explicitly incorporating legal soundness and structural completeness for reliable automated assessment.
Limitations
This study has several limitations, which suggest directions for future research. First, the current dataset is limited to a specific technical domain and time period. Expanding it to cover a broader range of technologies and filing dates could improve the generalizability of our framework. Second, PatentScore focuses on intrinsic aspects of claim quality,such as structural, legal, and semantic factors,but does not address extrinsic properties such as novelty or distinctiveness. Future work may incorporate comparative modules that evaluate overlap with prior art, enabling assessments of claim uniqueness and scope validity, which are critical in real-world patent examination. Third, although PatentScore is primarily designed for evaluating LLM-generated claims, its modular architecture allows potential adaptation to human-written patents. This could support automated analysis and feedback on structural coherence, clarity, and legal soundness in expert-authored claims. Finally, integrating legal knowledge graphs into the generation and evaluation pipeline could enhance factual accuracy and contextual alignment.
Ethical Considerations
In conducting this research, we ensured the highest ethical standards were upheld. All data used in this study, including GPT-4o-mini generated claims and the HUPD dataset, were obtained from publicly available sources and used in compliance with the respective licenses, such as the CC BY License. Human evaluators who participated in the evaluation process provided voluntary and informed consent before contributing to the study. Their identities remain anonymous to protect their privacy and confidentiality.
The study does not involve any personally identifiable information or sensitive data, and it strictly adheres to ethical guidelines for research involving human evaluators. Furthermore, we acknowledge the potential for misuse of AI in automating patent generation and emphasize the importance of developing AI tools that align with legal and ethical standards. The proposed framework, PatentScore, aims to improve transparency, reliability, and accountability in AI-generated patent claims.
Acknowledgments
We thank the creators of the HUPD dataset (CC BY License) and OpenAI for access to GPT-4o-mini, which were essential for this study. We also thank the three evaluators for their valuable contributions and voluntary participation after providing informed consent. Additionally, quantifying legal scope accuracy and completeness through metrics like Precision, Recall, and F-score could be investigated to assess claim quality against groundtruth data when a large labelled claim generation database is available.
Appendix A Appendix
A.1 Prompts for Experiments
A.1.1 Prompts for Claim Structure Evaluation
Name: Patent Claim Structure
I. Overview
Definition: A patent claim’s composition consisting of three essential parts (preamble, transitional phrase, and elements), which verify whether all these elements are properly positioned.
Evaluation Purpose: To systematically assess the structural completeness and proper organization of a patent claim to ensure clear and enforceable rights.
II. Key Evaluation Points
  1. 1.
    Preamble identification and content validation
  2. 2.
    Transitional phrase appropriateness and placement
III. Important Notes
  • Ensure all three essential parts exist.
  • Check proper element positioning.
IV. Standard Examples
"A device for processing data, comprising: a memory configured to store data; a processor coupled to the memory; and an interface connected to the processor."
V. Evaluation Procedure
  1. 1.
    Content Identification
    Input: Claim 1 requiring evaluation
    Output: Checklist, mapping, and element classification
    Q: Have all structural elements been identified?
    Tasks:
    • Carefully read the provided Claim 1 to identify:
      • Preamble
      • Transitional phrase
      • Elements/components
    • Verify presence of all essential parts:
      • Preamble completeness
      • Appropriate transitional phrase
      • All necessary elements
    • Create mapping of identified elements.
  2. 2.
    Critical Element Verification
    Input: Identified elements and relationships
    Output: Verification report of required elements
    Q: Do the essential elements meet requirements?
    Tasks:
    • Confirm essential structural elements:
      • Clear and complete preamble statement
      • Standard transitional phrase usage
      • Properly structured body elements
    • Verify element relationships:
      • Proper hierarchical structure
    • Remember missing or inappropriate items
  3. 3.
    Format Compliance Assessment
    Input: Element verification results
    Output: Compliance evaluation report
    Q: Does it comply with format requirements?
    Tasks:
    • Check format compliance:
      • Indentation and spacing
      • Punctuation usage (colons, semicolons)
      • Conjunction placement
    • Identify rule violations
    • Remember formatting errors
  4. 4.
    Standard-Based Comparative Analysis
    Input: Compliance assessment results
    Output: Comparative analysis report
    Q: How well does it match standards/examples?
    Tasks:
    • Compare with standard claim structures:
      • Similar technology field examples
      • Accepted structural patterns
    • Identify structural deviations
    • Remember errors
  5. 5.
    Final Scoring
    Input: Analysis results from previous steps
    Output: Final numerical score
    Patent Claim Structure Scoring (1-5 points):
    • 1 point (<20%):
      Example: "Processing data with memory and processor."
      Issues: No clear preamble, no transitional phrase, unstructured elements, missing punctuation, no formatting.
    • 2 points (20-40%):
      Example: "A data processing device having memory for storing data processor for processing interface for connecting."
      Issues: Weak preamble, missing transitional phrase, poor organization, incorrect punctuation, basic elements only.
    • 3 points (41-60%):
      Example: "A device for processing data comprising: a memory that stores data a processor connected to memory an interface; wherein the processor processes data."
      Issues: Basic preamble, has transitional phrase, basic formatting, some punctuation errors, basic relationships.
    • 4 points (61-80%):
      Example: "A device for processing data, comprising: a memory configured to store data; a processor coupled to the memory; an interface connected to the processor; wherein the interface transmits data."
      Strengths: Clear preamble, correct transitional phrase, good formatting, minor structural issues, clear relationships.
    • 5 points (>80%):
      Example: "A device for processing data, comprising: a memory configured to store input data; a processor coupled to the memory and configured to process data; an interface coupled to the processor; wherein the processor updates rules based on feedback."
      Strengths: Perfect structure, proper formatting, clear hierarchy, complete elements, professional composition.
A.1.2 Prompts for Punctuation in Patent Claims
Name: Punctuation in Patent Claim
I. Overview
Definition: Evaluation of proper use and placement of essential punctuation marks in patent claims (colon, semicolons, "and", period).
Evaluation Purpose: To systematically assess the accuracy of punctuation usage to ensure clarity and readability of the claim.
II. Key Evaluation Points
  1. 1.
    Essential punctuation mark validation
  2. 2.
    Appropriate punctuation placement
III. Important Notes
  • Check for colon (:) after transitional phrase.
  • Check for semicolons (;) after elements.
  • Check for "and" before the last element.
  • Check for period (.) at the end of claim.
IV. Standard Examples
"A device for processing data, comprising: a memory configured to store data; a processor coupled to the memory; and an interface connected to the processor."
V. Evaluation Procedure
  1. 1.
    Content Identification
    Input: Claim 1 requiring evaluation
    Output: Checklist, mapping, and punctuation classification
    Q: Have all punctuation marks been identified?
    Tasks:
    • Carefully read the provided Claim 1 to identify:
      • Colon location
      • Semicolon locations
      • "And" location
      • Period location
    • Verify presence of all essential punctuation:
      • Colon after transition
      • Semicolons after each element
      • "And" before last element
      • Period at claim end
    • Create mapping of identified punctuation
  2. 2.
    Critical Element Verification
    Input: Identified punctuation and locations
    Output: Essential punctuation verification report
    Q: Do the essential punctuation marks meet requirements?
    Tasks:
    • Confirm essential punctuation:
      • Appropriate colon usage
      • Appropriate semicolon usage
      • Appropriate "and" usage
    • Verify punctuation locations:
      • Exact position of each mark
    • Remember missing or inappropriate items
  3. 3.
    Format Compliance Assessment
    Input: Punctuation verification results
    Output: Compliance evaluation report
    Q: Does punctuation comply with format requirements?
    Tasks:
    • Check punctuation format compliance:
      • Spacing around punctuation
      • Consistency in usage
    • Remember formatting errors
  4. 4.
    Standard-Based Comparative Analysis
    Input: Compliance assessment results
    Output: Comparative analysis report
    Q: How well does it match standards/examples?
    Tasks:
    • Compare with standard punctuation usage:
      • Standard case patterns
      • Accepted usage practices
    • Identify punctuation deviations
    • Remember errors
  5. 5.
    Final Scoring
    Input: Analysis results from previous steps
    Output: Final numerical score
    Patent Claim Punctuation Scoring (1-5 points):
    • 1 point (<20%):
      Example: "A device with memory processor and interface."
      Issues: No colon after transition, no semicolons between elements, missing "and", missing period, no spacing after marks.
    • 2 points (20-40%):
      Example: "A device comprising memory; processor; interface."
      Issues: Missing colon after transition, inconsistent semicolon use, missing "and", has period, poor spacing.
    • 3 points (41-60%):
      Example: "A device comprising: memory; processor; and interface."
      Issues: Has colon, some semicolons, has "and", missing period, basic spacing.
    • 4 points (61-80%):
      Example: "A device, comprising: memory; processor; and interface."
      Strengths: Proper colon, most semicolons correct, proper "and", has period, minor spacing issues.
    • 5 points (>80%):
      Example: "A device for processing data, comprising: a memory configured to store data; a processor coupled to the memory; and an interface connected to the processor."
      Strengths: Perfect colon placement, all semicolons correct, proper "and" placement, correct period, perfect spacing.
A.1.3 Claim Inconsistent Element Referencing
Name: Claim Inconsistent Element Referencing
I. Overview
Definition: Assessment of consistency in referring to claim components (e.g., touchscreen display, processor, battery), ensuring each element maintains uniform terminology throughout the claim and avoids ambiguous references.
Evaluation Purpose: To ensure each component in the claim is consistently referenced using precise terminology, maintaining clarity and avoiding confusion in element relationships.
II. Key Evaluation Points
  1. 1.
    Terminology consistency verification
  2. 2.
    Reference clarity assessment
III. Important Notes
  • Check for consistent use of component terms
  • Verify clear reference relationships
  • Confirm no ambiguous references
IV. Standard Examples
"A device, comprising: a touchscreen display; a processor connected to the touchscreen display; and a battery powering the processor and the touchscreen display."
V. Evaluation Procedure
  1. 1.
    Content Identification
    Input: Claim 1 requiring evaluation
    Output: Checklist, mapping, and reference classification
    Q: Have all element references been identified?
    Tasks:
    • Carefully read the provided Claim 1 to identify:
      • Each component term
      • Each reference to components
      • Reference relationships
    • Verify component terminology:
      • Initial component definitions
      • Subsequent references
      • Reference consistency
    • Create mapping of identified references
  2. 2.
    Critical Element Verification
    Input: Identified references and relationships
    Output: Reference verification report
    Q: Do the reference terms meet requirements?
    Tasks:
    • Confirm reference consistency:
      • Uniform terminology use
      • Clear reference connections
      • Proper reference terms
    • Verify reference accuracy:
      • Correct component references
    • Remember inconsistent or improper items
  3. 3.
    Format Compliance Assessment
    Input: Reference verification results
    Output: Compliance evaluation report
    Q: Does reference usage comply with format requirements?
    Tasks:
    • Check reference format compliance:
      • Reference term consistency
      • Reference terminology clarity
    • Remember formatting errors
  4. 4.
    Standard-Based Comparative Analysis
    Input: Compliance assessment results
    Output: Comparative analysis report
    Q: How well does it match standards/examples?
    Tasks:
    • Compare with standard reference patterns:
      • Standard reference cases
      • Accepted reference practices
    • Identify reference deviations
    • Remember errors
  5. 5.
    Final Scoring
    Input: Analysis results from previous steps
    Output: Final numerical score
    Patent Claim Element Reference Consistency Scoring (1-5 points):
    • 1 point (<20%):
      Example: "A device including a screen connected to CPU, wherein the display is connected to processor, and a battery powering the touchscreen."  Issues: Different terms for same element (screen/display/touchscreen), incorrect reference terms (CPU/processor), unclear reference relationships, serious inconsistencies, confusing element relationships.
    • 2 points (20-40%):
      Example: "A touchscreen display; a processor connected to the screen; and a battery powering the CPU and display."
      Issues: Some inconsistent terminology, mixed reference terms, unclear reference relationships, major inconsistencies, basic element mentions only.
    • 3 points (41-60%):
      Example: "A touchscreen display; a processor connected to the touchscreen display; and a battery powering the processor and display."  Issues: Basic consistency, some missing references, partially clear relationships, some inconsistencies, basic reference structure.
    • 4 points (61-80%):
      Example: "A touchscreen display; a processor connected to said touchscreen display; and a battery powering said processor and said touchscreen display."
      Strengths: Mostly consistent terminology, proper reference terms, clear reference relationships, minor inconsistencies, clear element relationships.
    • 5 points (>80%):
      Example: "A device, comprising: a touchscreen display; a processor operatively connected to said touchscreen display; and a battery configured to power said processor and said touchscreen display."
      Strengths: Perfect terminology consistency, accurate reference terms, clear reference relationships, professional composition, perfect element relationships.
A.1.4 Patent Claim Antecedent Basis
Name: Patent Claim Antecedent Basis
I. Overview
Definition: Assessment of proper introduction and subsequent referencing of claim elements, ensuring each "the" reference has a corresponding previous introduction with "a" or "an".
Evaluation Purpose: To ensure clear and legally proper element referencing by confirming proper establishment of antecedent basis for all claim elements.
II. Key Evaluation Points
  1. 1.
    Article usage verification (a, an, the)
  2. 2.
    First mention validation
III. Important Notes
  • Check first introduction uses "a" or "an"
  • Verify subsequent references use "the"
  • Confirm all "the" references have prior introduction
IV. Standard Examples
"A device, comprising: a display; a processor connected to the display; and a battery, wherein the battery powers the processor and the display."
V. Evaluation Procedure
  1. 1.
    Content Identification
    Input: Claim 1 requiring evaluation
    Output: Checklist, mapping, and article usage classification
    Q: Have all article usages been identified?
    Tasks:
    • Carefully read the provided Claim 1 to identify:
      • First mentions (a, an)
      • Subsequent references (the)
      • Reference relationships
    • Verify article usage:
      • Initial introductions
      • Subsequent references
      • Reference consistency
    • Create mapping of identified articles.
  2. 2.
    Critical Element Verification
    Input: Identified articles and relationships
    Output: Article usage verification report
    Q: Do the article usages meet requirements?
    Tasks:
    • Confirm proper introduction:
      • Proper first mention articles
      • Proper subsequent articles
    • Verify reference accuracy:
      • Match between references
    • Remember improper article usage
  3. 3.
    Format Compliance Assessment
    Input: Article verification results
    Output: Compliance evaluation report
    Q: Does article usage comply with requirements?
    Tasks:
    • Check article usage compliance:
      • Article sequence
      • Reference clarity
    • Remember formatting errors
  4. 4.
    Standard-Based Comparative Analysis
    Input: Compliance assessment results
    Output: Comparative analysis report
    Q: How well does it match standards/examples?
    Tasks:
    • Compare with standard article usage:
      • Standard reference patterns
      • Accepted article practices
    • Identify article usage deviations
    • Remember errors
  5. 5.
    Final Scoring
    Input: Analysis results from previous steps
    Output: Final numerical score
    Patent Claim Antecedent Basis Scoring (1-5 points):
    • 1 point (<20%):
      Example: "The display and the CPU, wherein the memory connects to the processor."
      Issues: Improper first mentions, references without antecedent basis, incorrect article usage, severe antecedent basis issues, confusing relationships.
    • 2 points (20-40%):
      Example: "The display and a processor, wherein the CPU connects to memory."
      Issues: Some improper first mentions, incomplete antecedent basis, incorrect article usage, major antecedent basis issues, basic elements only.
    • 3 points (41-60%):
      Example: "A display; the processor connected to the display; and the memory connected to the processor."
      Issues: Basic antecedent basis, some incorrect articles, some unclear references, some antecedent basis issues, basic reference structure.
    • 4 points (61-80%):
      Example: "A display; a processor connected to the display; and the processor connected to a memory."
      Strengths: Mostly proper antecedent basis, proper article usage, clear references, minor antecedent basis issues, clear element relationships.
    • 5 points (>80%):
      Example: "A display; a processor connected to the display; a memory; and a battery connected to the processor and the memory."
      Strengths: Perfect antecedent basis, accurate article usage, clear references, professional composition, perfect element relationships.
A.1.5 Claim Validity & Uniqueness
Name: Claim Validity & Uniqueness
I. Overview
Definition: Evaluation of whether each claim component contributes new technical features and specific elements to the invention, without contradictions or unnecessary repetitions of earlier components.
Evaluation Purpose: To ensure the claim maintains its validity and uniqueness by differentiating each part and upholding legal strength within the patent framework.
II. Key Evaluation Points
  1. 1.
    Technical feature uniqueness
  2. 2.
    Component differentiation
III. Important Notes
  • Verify each component adds unique value
  • Check for technical contradictions
  • Confirm no unnecessary repetitions
IV. Standard Examples
"A data processing device, comprising: a unique processing unit having specific computational capabilities; a specialized memory unit configured to enhance processing speed; and a novel interface designed to optimize data transfer."
V. Evaluation Procedure
  1. 1.
    Content Identification
    Input: Claim 1 requiring evaluation
    Output: Checklist, mapping, and uniqueness classification
    Q: Have all unique elements been identified?
    Tasks:
    • Carefully read the provided Claim 1 to identify:
      • Technical features
      • Unique elements
      • Specific contributions
    • Verify element uniqueness:
      • Technical differentiation
      • Specific features
    • Create mapping of unique features
  2. 2.
    Critical Element Verification
    Input: Identified elements and features
    Output: Validity verification report
    Q: Do the elements contribute unique value?
    Tasks:
    • Confirm technical contribution:
      • Novel features
      • Specific improvements
    • Verify element differentiation:
      • Technical distinctions
    • Remember contradictions or repetitions
  3. 3.
    Format Compliance Assessment
    Input: Verification results
    Output: Compliance evaluation report
    Q: Does the claim maintain validity requirements?
    Tasks:
    • Check validity compliance:
      • Technical consistency
      • Feature uniqueness
    • Remember validity issues
  4. 4.
    Standard-Based Comparative Analysis
    Input: Compliance assessment results
    Output: Comparative analysis report
    Q: How well does it establish uniqueness?
    Tasks:
    • Compare with standard features:
      • Similar technology features
      • Distinctive elements
    • Identify uniqueness deviations
    • Remember errors.
  5. 5.
    Final Scoring
    Input: Analysis results from previous steps
    Output: Final numerical score
    Patent Claim Validity & Uniqueness Scoring (1-5 points):
    • 1 point (<20%):
      Example: "A data processing device comprising a processor and memory."
      Issues: No technical features, generic components only, lack of differentiation, no uniqueness, vague implementation.
    • 2 points (20-40%):
      Example: "A data processing device with a high-speed processor and large memory."
      Issues: Minimal technical features, basic differentiation attempt, insufficient specificity, limited uniqueness, common technical level.
    • 3 points (41-60%):
      Example: "A data processing device comprising: a parallel processing processor; a high-speed cache memory; and a data optimization interface."
      Issues: Basic technical features, some differentiation, basic uniqueness, partial specificity, some technical improvements.
    • 4 points (61-80%):
      Example: "A data processing device comprising: a multi-core parallel processor; a hierarchical cache memory structure; and a real-time data optimization interface."
      Strengths: Clear technical features, specific differentiation, good uniqueness, concrete implementation, clear technical improvements.
    • 5 points (>80%):
      Example: "A data processing device comprising: an AI-accelerated multi-core processor; a dynamic memory allocation hierarchical cache; and an intelligent interface with adaptive data optimization algorithms."
      Strengths: Strong technical features, excellent differentiation, high uniqueness, detailed implementation, innovative improvements.
A.1.6 Ambiguous Claim Scope
Name: Ambiguous Claim Scope
I. Overview
Definition: Evaluation of claim language clarity and boundary definition, ensuring specific and unambiguous protection scope without unclear or overly broad terms that could lead to uncertainty in claim interpretation.
Evaluation Purpose: To ensure the claim clearly defines its protection boundaries using specific and precise language, avoiding ambiguous or overly broad terms that could weaken legal protection.
II. Key Evaluation Points
  1. 1.
    Boundary clarity verification
  2. 2.
    Term specificity assessment
III. Important Notes
  • Check for clear scope boundaries
  • Verify specific technical terms
  • Confirm no ambiguous language
IV. Standard Examples
"A data processing device, comprising: a 2.4 GHz processor specifically configured for image processing; a 512GB solid-state memory connected to the processor; and a 4K resolution display interface operating at 60Hz refresh rate."
V. Evaluation Procedure
  1. 1.
    Content Identification
    Input: Claim 1 requiring evaluation
    Output: Checklist, mapping, and scope classification
    Q: Have all scope-defining elements been identified?
    Tasks:
    • Carefully read the provided Claim 1 to identify:
      • Scope boundaries
      • Technical specifications
      • Limiting terms
    • Verify term specificity:
      • Technical parameters
      • Functional limitations
    • Create mapping of scope elements
  2. 2.
    Critical Element Verification
    Input: Identified scope elements and boundaries
    Output: Scope verification report
    Q: Do the elements clearly define boundaries?
    Tasks:
    • Confirm boundary clarity:
      • Clear technical limits
      • Specific definitions
    • Verify scope precision:
      • Term specificity
    • Remember ambiguous or broad terms
  3. 3.
    Format Compliance Assessment
    Input: Scope verification results
    Output: Compliance evaluation report
    Q: Does the scope comply with clarity requirements?
    Tasks:
    • Check scope compliance:
      • Term precision
      • Boundary clarity
    • Remember clarity issues
  4. 4.
    Standard-Based Comparative Analysis
    Input: Compliance assessment results
    Output: Comparative analysis report
    Q: How well does it define protection scope?
    Tasks:
    • Compare with standard scope definitions:
      • Similar technology scopes
      • Accepted boundary definitions
    • Identify scope deviations
    • Remember errors
  5. 5.
    Final Scoring
    Input: Analysis results from previous steps
    Output: Final numerical score
    Patent Claim Scope Ambiguity Scoring (1-5 points):
    • 1 point (<20%):
      Example: "An improved device that processes data better."
      Issues: Highly ambiguous terms, unclear boundaries, unmeasurable improvements, subjective expressions, boundless protection scope.
    • 2 points (20-40%):
      Example: "An enhanced processing device with a high-speed processor and large memory."
      Issues: Vague technical terms, broad scope, unclear parameters, generic expressions, wide protection scope.
    • 3 points (41-60%):
      Example: "A processing device comprising: a processor operating above 1GHz; and memory exceeding 256GB."
      Issues: Basic clarity, some specific parameters, basic boundary definition, measurable features, limited protection scope.
    • 4 points (61-80%):
      Example: "A processing device comprising: a 2.4GHz multi-core processor; a 512GB SSD; and a 4K resolution display."
      Strengths: Clear technical terms, specific parameters, clear boundaries, precise measurements, defined protection scope.
    • 5 points (>80%):
      Example: "A processing device comprising: a 2.4GHz 8-core processor performing 1M operations/second; a 512GB NVMe SSD with 3000MB/s read speed; and a 4K 60Hz HDR display."
      Strengths: Very precise technical terms, clear performance parameters, exact boundary definitions, specific measurement criteria, precise protection scope.
Appendix B Prompts for generating by LLM
You are an experienced patent attorney with a deep understanding of patent law and intellectual property (IP) protection. I will provide you with the title, abstract, and detailed description of a specific invention. Based on this information, please draft Claim 1 of the patent, ensuring that it accurately captures the essence of the invention while being legally strong and precise. Claim 1 should encompass the core technology of the invention while meeting legal requirements to distinguish it from prior art. It should use a clear and specific terminology to define the scope of protection appropriately, avoiding overly narrow or unnecessarily broad expressions.
Additionally, the structure of Claim 1 should adhere to the conventional format used in patent legal documents and incorporate terminology commonly used in the relevant technical field.
The Claim 1 you draft must meet the following criteria:
1. Clarity: Avoid ambiguous expressions and use precise language that can be easily understood. 2. Legal Validity: Comply with patent examination standards and be legally appropriate to provide substantial protection. 3. Breadth & Differentiation: Adequately cover the core elements of the invention while highlighting its distinction from existing technology. 4. Technical Accuracy: Accurately reflect the provided invention description without technical errors. 5. Structural Integrity: Follow proper grammar and format for patent claims and maintain logical consistency.
Now, I will provide the patent title, abstract, and detailed description. Based on this information, please draft Claim 1 of the patent.