### v2

# PatentScore: Multi-dimensional Evaluation of LLM-Generated Patent Claims

Abstract
High-stakes texts such as patent claims, medical records, and technical reports are structurally complex and demand a high degree of reliability and precision. While large language models (LLMs) have recently been applied to automate their generation in high-stakes domains, reliably evaluating such outputs remains a major challenge. Conventional natural language generation (NLG) metrics are effective for generic documents but fail to capture the structural and legal characteristics essential to evaluating complex high-stakes documents. To address this gap, we propose PatentScore, a multi-dimensional evaluation framework specifically designed for one of the most intricate and rigorous domains, patent claims. PatentScore integrates hierarchical decomposition of claim elements, validation patterns grounded in legal and technical standards, and scoring across structural, semantic, and legal dimensions. In experiments on our dataset which consists of 400 Claim1, PatentScore achieved the highest correlation with expert annotations ($r=0.819$), significantly outperforming widely used NLG metrics. This work establishes a new standard for evaluating LLM-generated patent claims, providing a solid foundation for research on patent generation and validation.
1 Introduction
Large language models (LLMs) have demonstrated impressive capabilities in automated text generation across a wide range of natural language processing tasks. However, existing evaluation frameworks remain insufficient for rigorously and systematically assessing their outputs in high-stakes domains. These domains, such as patent claims, legal contracts, and technical reports, involve deeply hierarchical and interdependent structures that cannot be accurately evaluated using surface-level metrics such as sentence-level fluency or shallow contextual similarity. Current evaluation metrics often fail to capture critical domain-specific nuances, including legal constraints, structural dependencies, technical precision, and implicit reasoning, which significantly limit their effectiveness in these contexts. This limitation is especially consequential because it poses risks to both legal validity and practical utility of the generated documents, thereby hindering their adoption in real-world, risk-sensitive applications (Guha et al., 2023).
Figure 1: A comparison of PatentScore and standard NLG metrics in evaluating structural and legal accuracy of patent claims.
Among high-stakes legal and technical documents, patent claims pose particularly significant challenges for evaluation. Because they function as both technical descriptions and legally binding texts, patent claims require precise assessment of structural elements such as antecedent consistency and claim dependency. Existing natural language generation (NLG) evaluation methods struggle in domains where both linguistic precision and legal accountability are critical. Patent claims exemplify these characteristics, as they define the scope of rights granted to an invention. The first claim, in particular, plays a central role in shaping the structure and coverage of the entire claim set. This pivotal role entails substantial socio-economic implications, as it determines the legal and commercial value of the invention (Merges, 1997; Bessen and Meurer, 2009; Surden, 2019).
With the rapid advancement of LLMs, recent studies have explored their potential to generate and evaluate precision-critical legal texts. However, patent claims remain a particularly challenging domain due to their intricate structural and legal constraints (Aristodemou and Tietze, 2018; Lee and Hsiang, 2020). Traditional NLG metrics, such as BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and BERTScore (Zhang et al., 2020), primarily focus on lexical overlap or contextual similarity, failing to address the structural and legal nuances required for patent claim evaluation.
More recent LLM-based evaluation approaches, such as LLM-as-a-Judge (Zheng et al., 2023), rely on general-purpose models to assess fluency and semantic coherence. However, they lack the domain-specific knowledge required to evaluate structural constraints, including claim dependency and legal clarity, that are critical to patent claims. These shortcomings limit the applicability of such metrics in high-stakes legal contexts where precision is paramount (Zuo et al., 2024). For instance, inconsistent antecedent references can render a patent legally ambiguous, while punctuation errors may distort its scope of protection (Merges, 1997).
To address these limitations, we propose PatentScore, a multi-dimensional evaluation framework for patent claims. This framework evaluates structural and legal elements such as antecedent consistency, claim dependency, and legal clarity. Our approach decomposes claims into constituent parts and applies targeted metrics, enabling reliable and consistent assessment of LLM-generated claims. This systematic evaluation significantly enhances the quality of claim evaluation compared to existing metrics. Our objective is to develop a domain-specific metric for patent claims while establishing a generalizable, structure-aware evaluation framework applicable to high-precision documents, such as contract clauses, policy reports, and regulatory compliance in clinical documentation.
The key contributions of PatentScore are summarized as follows:
  • Claim-Structured Evaluation Framework. We introduce an architecture that decomposes a claim into structural components and quantifies each, covering claim structure, element linkage, antecedent basis, and claim dependency per official examination guidelines.
  • Expert-Aligned Validation. We compare PatentScore’s evaluation of LLM-generated claims against expert ratings, achieving a strong correlation ($r=0.819$) and outperforming standard NLG metrics in capturing structural and legal precision.
  • Open Evaluation Benchmark. We release 400 LLM-generated claims with expert annotations, creating the first public benchmark for evaluating LLM-generated patent texts.
Although our experiments focus on patent claims, PatentScore is model-agnostic. Its structural and legal principles extend to other high-precision texts such as contract clauses, policy reports, clinical documentation, and regulatory filings. By providing a robust framework for evaluating structured texts, our work lays the foundation for establishing reliable evaluation protocols essential for the responsible use of LLMs in high-stakes domains.
2 Background and Related Work
2.1 Specifications in Patent Claims
Claim 1 (Example)
1. An extendible trailer tongue comprising:
  a first tubular member;
  a second tubular member telescopically received within the first tubular member; and
  a locking mechanism to secure the second tubular member in a selected position relative to the first tubular member.
Table 1: A Brief summary example of Claim 1 from the patent titled “Multiple component headgear system” (Siprut, 2000).
Patents are specialized legal-technical documents designed to protect intellectual property (IP) rights. Among their components, claims play a central role, as they define the precise legal boundaries of protection granted to an invention. In particular, Claim 1, which is the first claim, typically serves as the cornerstone of the claim set. It establishes the broadest scope of protection and strongly influences both the validity and enforceability of the patent as a whole (World Intellectual Property Organization, 2022; Merges, 1997). The breadth of Claim 1 not only constrains subsequent dependent claims but also directly affects the legal and commercial value of the invention.
Unlike general technical documents, patent claims must strictly adhere to structural, linguistic, and legal requirements. They follow a standardized format mandated by patent offices, employ precise domain-specific terminology, and are drafted with enforceability in mind. Table 1 illustrates an example of a well-structured Claim 1, which demonstrates these key requirements:
  • Structural format: Compliance with strict grammatical and formatting conventions, including numbered clauses, transitional phrases (e.g., “comprising”), and hierarchical listing of elements. Deviations can render a claim unclear or legally unenforceable.
  • Technical clarity: Precise definition of technical element and its relationships, ensuring that the scope of the invention is unambiguous to examiners, practitioners, and courts.
  • Terminology consistency: Reliance on standardized patent-specific language and consistent referencing of terms, minimizing ambiguity and potential disputes.
Together, these interrelated requirements ensure that Claim 1 functions simultaneously as a detailed technical specification and as a legally binding definition of the scope of the invention. Drafting Claim 1, therefore, requires a delicate and deliberate balance: it must be broad enough to secure meaningful protection and commercial value, while being sufficiently precise and consistent to withstand rigorous legal scrutiny.
2.2 Patent Generation and Evaluation
Automating patent generation and evaluation has recently emerged as an important research agenda, yet it remains relatively underexplored. LLMs have been increasingly applied to generate technical and legal content (Brown et al., 2020), showing promise in zero-shot reasoning (Kojima et al., 2022) and complex instruction following (Lai et al., 2024). However, most existing work has focused on general text generation, without addressing the strict structural and legal constraints unique to patent claims (Surden, 2019).
Patent evaluation inherently involves both linguistic and domain-specific dimensions. Traditionally, examiners and attorneys manually review claims, particularly Claim 1, which is labor-intensive and often inconsistent. With the rapid global increase in patent applications, automated evaluation systems must be able to simultaneously assess technical correctness, legal soundness, and structural compliance (Surden, 2019).
Nevertheless, evaluating LLM-generated patent claims remains a highly challenging and technically demanding task. Even minor errors in structure, terminology, or legal scope can result in costly litigation or eventual invalidation of the patent. Existing NLG metrics tend to emphasize surface-level linguistic similarity or superficial fluency, but they fail to capture critical patent-specific aspects such as claim scope, legal consistency, structural validity, and the intricate interdependence of claim elements. This limitation strongly underscores the necessity for domain-specific evaluation frameworks that can comprehensively and reliably address the unique requirements of patent claims.
2.3 NLG Metrics for Patent Evaluation
Traditionally, the evaluation of patent texts has relied on metrics using n-grams such as BLEU and ROUGE, which measure lexical overlap between generated and reference texts. However, these approaches overlook semantic relationships and domain-specific nuances, and thus are insufficient for patent claims that demand high levels of legal and technical precision. With the introduction of Transformer-based models, more nuanced evaluation methods became possible. For example, BERTScore leverages contextual embeddings to assess semantic similarity (Zhang et al., 2020), while BLEURT combines BERT’s representational power with task-specific optimization (Sellam et al., 2020).
More recently, LLM-based evaluation frameworks have emerged, directly employing large language models as evaluators. GPTScore exploits the reasoning capability of LLMs to assess semantic coherence, logical consistency, and contextual appropriateness (Fu et al., 2024). Other studies further highlight the potential of LLM-as-a-Judge (Zheng et al., 2023), showing that LLM evaluators can often align more closely with human judgment compared to traditional automatic metrics.
However, these approaches still fail to adequately capture the strict legal and structural constraints inherent in patent claims. Critical aspects such as claim dependency, antecedent basis, and legal clarity remain largely beyond the practical scope of existing evaluation metrics. Unlike general text, patent claims require rigorous structural compliance, precise legal terminology, and systematic domain-specific validation. This persistent gap underscores the urgent need for specialized evaluation frameworks that can comprehensively address both linguistic quality and patent-specific requirements. PatentScore is explicitly designed to fill this void by holistically integrating structural, semantic, and legal dimensions into a unified framework tailored for patent claims.
Figure 2: PatentScore Evaluation Framework: decomposition into structural, legal, and semantic dimensions, normalization of metrics, and weighted aggregation for final scoring.
3 PatentScore for Patent Claim Evaluation
3.1 Evaluation Dimensions of PatentScore
PatentScore is a multi-dimensional evaluation framework that assesses the quality of patent claims by decomposing them into three independent aspects: structural, legal, and semantic. Each dimension is grounded in domain-specific standards, particularly the WIPO Patent Drafting Manual World Intellectual Property Organization (2022) and the USPTO USPTO (2023), and is operationalized through distinct component metrics. These dimensions operate independently, allowing for a detailed diagnosis of various characteristics of a patent claim.
(a) Structural Dimension.
The structural dimension evaluates the syntactic organization and formal composition of patent claims. It focuses on whether the claim conforms to internationally accepted formatting rules that promote legal clarity and technical completeness. This dimension comprises the following three components:
  • Claim Structure ($M_{CS}$): Checks whether the claim follows the canonical structure, including a preamble, constituent elements, and functional linkages, and logical coherence between interconnected parts.
  • Claim Punctuation ($M_{CP}$): Verifies proper punctuation use (e.g., commas, semicolons) that affect precise clause separation and legal interpretation.
  • Antecedent Basis ($M_{AB}$): Assesses the proper usage of definite noun phrases (e.g., “the module”, “said component”) that must be supported by clear prior mention.
(b) Legal Dimension.
The legal dimension assesses whether the claim satisfies enforceability criteria under patent law, referencing legal doctrine from the USPTO’s MPEP and WIPO standards. Its core metrics include:
  • Element Referencing ($M_{ER}$): Measures proper dependency and referencing in independent and dependent claims.
  • Validity and Uniqueness ($M_{VU}$): Evaluates novelty, non-triviality, and absence of internal contradiction.
  • Ambiguous Scope ($M_{AS}$): Identifies vague or overbroad terms that could undermine legal clarity or enforceability.
(c) Semantic Dimension.
The semantic dimension evaluates the extent to which a generated claim preserves the intended meaning and technical context of the original. We adopt BERTScore ($M_{BS}$), which measures semantic similarity using contextual embeddings, due to its robustness in capturing meaning beyond lexical overlap. To ensure compatibility with other dimensions, BERTScore values in $[0,1]$ are linearly rescaled to a $[1,5]$ range.
This decomposition enables fine-grained diagnostics of claim quality, enhancing both the reliability and interpretability of LLM-based assessments.
3.2 Implementation of PatentScore
To operationalize the three evaluation dimensions, PatentScore employs an LLM-based prompt evaluation scheme. For each metric, a dedicated prompt is designed to guide the model’s attention to the corresponding evaluation criterion. GPT-4o-mini is used as the primary evaluator, and ten independent evaluations are conducted for each metric, with the average score adopted as the final result. This design reduces output variance and ensures stability in scoring.
Different evaluation strategies are applied depending on the nature of each metric. For structural and legal metrics, explicit prompts are used to focus the model on the core evaluation concepts. In contrast, the semantic metric is assessed using BERTScore, which is computed independently. The original $[0,1]$ range of BERTScore values is linearly rescaled to a $[1,5]$ range for consistency with other scores. All prompt templates are provided in detail in Appendix A.
The proposed implementation is model-agnostic and can be applied to various LLMs, including GPT-4o-mini and others. Since the evaluation criteria are grounded in internationally recognized patent drafting guidelines (WIPO, USPTO), the framework ensures generalizability and applicability across different contexts.
3.3 Prompt Design for Claim Evaluation
In this study, we propose an LLM-adaptable prompt framework for systematic evaluation of patent claims. Our prompt engineering methodology builds upon Chain of Thought (Wei et al., 2022) reasoning, extending its sequential inference by incorporating prompts that explicitly address structural, legal, and semantic requirements of patent claims. For instance, prompts for structural metrics guide the model to verify claim formatting, while legal prompts focus on compliance with patentability criteria. This expanded prompt structure ensures consistency across evaluations while enabling parallel processing of multiple evaluation dimensions, allowing for a more sophisticated and systematic analysis. Detailed prompts for each evaluation component are provided in Appendix A.
Component-specific Template Design We designed a standardized base template inspired by Chain of Thought, extending its sequential reasoning to a multidimensional framework tailored for patent claim evaluation, enabling both consistency and comprehensive analysis.
[…]
FOCUS: Component-specific evaluation
OUTPUT FORMAT: Detailed analysis and score (1-5)
[…]
Primary Requirements:
- [Component-specific requirement 1]
- [Component-specific requirement 2]
Secondary Requirements:
- [Additional evaluation points]
[…]
1. Identify relevant claim elements
2. Evaluate against requirements
3. Assign score with justification
[…]
5: Exceeds all requirements with exceptional clarity
4: Meets all requirements effectively
3: Meets primary requirements adequately
2: Partially meets requirements with deficiencies
1: Fails to meet critical requirements
Implementation by Components We customize the templates for structural analysis ($M_{CS}$, $M_{CP}$, $M_{AB}$) and legal compliance ($M_{ER}$, $M_{VU}$, $M_{AS}$) components. For semantic evaluation ($M_{BS}$), we utilize BERTScore, which provides an automated assessment of semantic similarity between generated and reference claims through contextual embeddings. This structured approach, combining specialized prompts for structural and legal assessments with automated semantic evaluation, enables a more reliable and fully reproducible evaluation of patent claim quality across all dimensions of our framework.
3.4 Score Integration and Weighting
PatentScore integrates seven metric scores into a final evaluation score by assigning a weight to each component. To determine the weights, we conduct ablation studies that quantify the relative importance of each evaluation metric.
For each component $M_{i}$, we compute its impact on performance by removing it from the evaluation and observing the changes in Pearson correlation with expert scores ($\Delta r_{i}=r_{\text{full}}-r_{i}$) and mean absolute error ($\Delta\text{MAE}_{i}=\text{MAE}_{i}-\text{MAE}_{\text{full}}$) relative to the full model. Specifically:
$\Delta\text{MAE}_{i}=\text{MAE}_{i}-\text{MAE}_{\text{full}},$ (1)
where $\text{MAE}_{\text{full}}$ are the performance metrics using all components, and $\text{MAE}_{i}$ are the metrics after removing component $M_{i}$. The final weight $w_{i}$ for component $M_{i}$ is then calculated as:
$w_{i}=\frac{\Delta r_{i}+\Delta\text{MAE}_{i}}{\sum_{j=1}^{7}(\Delta r_{j}+\Delta\text{MAE}_{j})}.$ (2)
To ensure compatibility, each metric score $M_{i}$ is normalized to the range $[1,5]$ before aggregation. The final PatentScore is computed as the weighted sum of all component scores:
$\text{PatentScore}=\sum_{i=1}^{7}w_{i}\cdot M_{i}.$ (3)
This weighted integration reflects the empirically derived contribution of each component and enhances the alignment of PatentScore with expert human evaluations. It rewards dimensions with higher diagnostic power to exert more influence on the final score, while preserving interpretability and fairness across evaluation aspects.
4 Data and Verification Measures
4.1 Claim Data Generation and Composition
Patent Number: US20170123456
Gold-Claim A method for processing biometric data, comprising: receiving sensor data; extracting features; and authenticating a user.
LLM-Claim A method for biometric processing, comprising: obtaining sensor data; processing features; and performing authentication.
Expert Scores E1: 3.0  E2: 4.0  E3: 3.0 Mean Score: 3.33
Table 2: An example of a human-authored patent claim (Gold-Claim) and its GPT-4o-mini-generated (LLM-Claim) counterpart, with expert annotations.
We use a subset of the HUPD dataset (Suzgun et al., 2022), consisting of patents filed in 2016 and 2017, selected for their recency and relevance to modern technological domains. This subset includes the first claim of 400 patents classified under Section A (Human Necessities) and Section G (Physics) of the International Patent Classification (IPC), with 200 claims from each section, ensuring a balanced representation of diverse technical fields. The dataset was curated to support the comparative evaluation of patent claims generated by GPT-4o-mini and those written by human experts. Table 2 presents an example entry, which serves as the basis for assessing the legal and technical completeness of generated claims, and for validating our proposed sequential evaluation metric. Each entry includes the patent ID, the original human-authored Claim 1, and the corresponding GPT-4o-mini-generated version. The dataset and implementation used in this study are publicly available at https://github.com/Yongmin-Yoo/PatentScore.
4.2 Expert Evaluation
To verify the reliability of LLM-generated claims and validate the effectiveness of PatentScore, we organized an evaluation panel comprising three domain experts: one legal expert specializing in patent law and two technical experts holding advanced degrees in electrical engineering and computer science, respectively. Each expert independently assessed the generated and reference claims using a 5-point Likert scale (1 = poor, 5 = excellent). The assessment was conducted under blind conditions to prevent bias, and each claim was evaluated on multiple criteria, including (i) structural compliance with patent claim conventions (e.g., preamble-body format, proper use of antecedents), (ii) clarity and precision in legal phrasing, and (iii) completeness and correctness of the technical disclosure.
To ensure reliability and reproducibility, we recorded all individual expert scores (E1–E3) and computed their mean for each claim pair, as summarized in Table 2. This multi-expert annotation provides a robust ground truth for benchmarking automatic evaluators and also enables error analysis in cases of inter-rater disagreement. The use of experts from both legal and technical domains ensures that PatentScore is validated against the multi-faceted nature of patent quality, thereby reinforcing its credibility as an evaluation framework.
4.3 Verification Measures
To validate the effectiveness of PatentScore, we conducted a comparative evaluation against widely used NLG metrics, including BLEU, ROUGE-L, BERTScore, GPTScore, and Cosine Similarity. As shown in Table 3, we assessed alignment with expert judgments using four statistical indicators: Pearson’s $r$, Spearman’s $\rho$, Kendall’s $\tau$, and mean absolute error (MAE). These quantitative measures capture correlation, ranking consistency, and deviation from expert scores, providing a comprehensive basis for evaluating the reliability of PatentScore relative to conventional metrics.
5 Experiment Results
5.1 Claim Generation Quality Analysis
Metric $r$ $\rho$ $\tau$ MAE
PatentScore .819 .813 .665 .568
PatentScore Avg. .818 .793 .551 .568
BERTScore -.161 -.163 -.130 1.975
GPT Score .013 -.004 -.003 1.408
BLEU -.117 -.089 -.065 1.744
ROUGE-L -.159 -.112 -.080 2.499
Cosine Sim -.050 .030 .024 1.946
Table 3: The comparison of various metrics using Pearson correlation $r$, Spearman correlation $\rho$, Kendall’s Tau $\tau$ and MAE.
Our framework demonstrates strong alignment with human expert evaluation, achieving a Pearson correlation coefficient of $r=0.819$ ($p<0.01$) across all metrics, substantially surpassing existing methods. It is noteworthy that the consistent negative correlation shown by traditional metrics such as BERTScore (r=0.161r=-0.161), ROUGE-L (r=0.159r=-0.159), and BLEU (r=0.117r=-0.117), indicating their limitations in patent claim evaluation. The effectiveness of the framework is further validated by its strong performance across multiple correlation measures, including Spearman’s rank correlation ($\rho=0.813$) and Kendall’s rank correlation ($\tau=0.665$). Beyond correlation coefficients, our approach demonstrates significantly lower error rates ($MAE=0.568$) compared to existing metrics, which exhibit substantially higher error values (e.g., $MAE=2.499$ for ROUGE-L).
Our results reveal that commonly used metrics show limited effectiveness in capturing the nuances of patent claim quality, as indicated by their negative correlations with expert judgments. While metrics like BERTScore and ROUGE-L show negative or near-zero correlation with expert scores, PatentScore exhibits strong positive correlation, indicating its suitability for the legal and structural nuances of patent claims. The consistently superior performance of our framework across diverse criteria establishes it as a robust and reliable method for patent claim assessment. To assess the generalizability of PatentScore beyond a single generation model, we conducted additional evaluations using GPT-generated claims from two distinct LLMs, Claude-3.5-Haiku and Gemini-1.5-Flash on the same 400-pair dataset under identical evaluation conditions. PatentScore achieved Pearson correlation coefficients of 0.745 and 0.731 on these models, respectively, indicating consistently strong alignment with expert annotations across diverse LLMs. These results reinforce the robustness and cross-model applicability of PatentScore as a reliable evaluator for patent claim quality.
5.2 Quality Comparison of Expert Evaluation on Generated Claims
The Pearson correlation coefficients between the ratings by three experts range from 0.795 to 0.847, demonstrating their high correlation and indicating that the evaluators followed similar standards in assessing claim quality. Cronbach’s $\alpha$ is 0.931, reflecting a very high level of internal consistency in their judgments across different claims. Furthermore, the ICC(3,k) value of 0.928 confirms excellent inter-rater agreement, while Krippendorff’s $\alpha$ of 0.784 suggests a substantial level of reliability even under more conservative assumptions. Taken together, these findings statistically validate that the proposed evaluation framework captures rigorous and consistent criteria, ensuring high levels of reliability and reproducibility in expert-based assessments. Importantly, such strong agreement among evaluators provides a solid foundation for benchmarking automated metrics, as it establishes that human judgments themselves are both stable and trustworthy, thereby reinforcing the credibility of subsequent comparisons between expert ratings and model-generated scores.
5.3 Ablation Study
Model Corr. ($r$) MAE Perf. Drop (%) Weight
Baseline ($\ddagger$) 0.818 0.568 - 1.0
- $M_{CS}$ 0.735 0.675 10.1 (18.8) 0.166
- $M_{CP}$ 0.723 0.667 11.6 (17.4) 0.171
- $M_{AB}$ 0.731 0.671 10.6 (18.1) 0.167
- $M_{ER}$ 0.734 0.670 10.3 (18.0) 0.163
- $M_{VU}$ 0.742 0.674 9.3 (18.7) 0.159
- $M_{AS}$ 0.729 0.677 10.9 (19.2) 0.173
- $M_{BS}$ 0.817 0.569 0.1 (0.2) 0.001
Notes: Numbers in parentheses indicate MAE increases. The baseline model ($\ddagger$) uses equal weights for all components.
Table 4: The comparison of metric contributions via ablation analysis.
We conducted an ablation study to analyze the impact of each metric on the overall framework performance, as shown in Table 4. The baseline model, which assigns equal weights (i.e., $w_{i}=1/7$) to all seven components, achieves a correlation of 0.818 with MAE of 0.568. Further, component-wise analysis reveals:
  • Structural metrics: The removal of structural components, including claim structure ($M_{CS}$), punctuation pattern ($M_{CP}$), and antecedent basis ($M_{AB}$), led to correlation drops of 10.1%, 11.6%, and 10.6%, respectively. These substantial decreases highlight the central role of structural compliance in patent claim evaluation. In particular, the antecedent basis metric, which ensures logical referential consistency across claim elements, emerges as a key determinant of expert-perceived quality.
  • Legal metrics: Excluding legal criteria, element referencing ($M_{ER}$), validity and uniqueness ($M_{VU}$), and ambiguous scope ($M_{AS}$), resulted in performance reductions of 10.3%, 9.3%, and 10.9%. These findings indicate that legal soundness is not a marginal factor but a core dimension in assessing claim quality. The ambiguous scope metric shows that evaluators strongly penalize imprecise or overly broad language, reaffirming the necessity of explicit legal rigor in automated evaluation.
  • Semantic metric: In contrast, the exclusion of the semantic metric BERTScore ($M_{BS}$) produced only a negligible decline of 0.1%. This suggests that while semantic similarity is not independently decisive, it provides complementary value by reinforcing lexical and terminological consistency, particularly in contexts involving emerging or domain-specific technical vocabulary.
Following the final weights in Eqn. 4, we can calculate the patent score, where each weight $W$ reflects the relative importance of corresponding metrics derived from the ablation study.
$\begin{split}\textit{PatentScore}=&\,0.166\cdot M_{CS}+0.171\cdot M_{CP}+\\ &\,0.163\cdot M_{ER}+0.167\cdot M_{AB}+\\ &\,0.159\cdot M_{VU}+0.173\cdot M_{AS}+\\ &\,0.001\cdot M_{BS}\end{split}$ (4)
These findings highlight the minimal impact of BERTScore on overall framework performance. Removing BERTScore results in only a 0.1% decrease, indicating that semantic metrics play a complementary rather than critical role in patent claim evaluation. In contrast, structural and legal metrics demonstrate substantial influence, with performance drops of 9.3% to 11.6% when excluded underscoring the importance of structural validity and legal compliance in determining claim quality. While modern LLMs exhibit strong semantic consistency and accuracy, they still require explicit guidance to meet structural and legal requirements. This balance of semantic strength and structural/legal support affirms the necessity of a comprehensive evaluation framework like PatentScore.
5.4 Key Findings
Limitations of Existing Metrics: Conventional text evaluation metrics (BLEU, ROUGE-L, and BERTScore) exhibit negative correlations ($r=-0.117\sim-0.161$) when applied to patent claim evaluation, underscoring their inadequacy for this specialized domain. BLEU’s reliance on n-gram overlap fails to capture logical dependencies, and ROUGE-L prioritizes surface-level similarity at the expense of legal and technical coherence. These limitations clearly highlight the need for a domain-tailored evaluation framework.
Balanced Component Evaluation: Our ablation study reveals comparable weights for structural (0.1660.1710.166-0.171) and legal components (0.1590.1730.159-0.173), demonstrating that both dimensions contribute equally to robust patent claim assessment. The 10.6% performance drop from omitting the antecedent basis metric ($M_{AB}$) illustrates how critical cross-referential clarity is in expert judgments. This finding confirms that a multi-faceted evaluation approach, rather than reliance on a single dimension, is essential for capturing claim quality.
Semantic Analysis Role: Although BERTScore contributes only marginal independent impact (weight: $0.001$), its inclusion plays a complementary role by reinforcing domain-specific terminology consistency, particularly for emerging technology terms. This demonstrates that semantic alignment alone is insufficient but still valuable as a supporting layer. Together with structural and legal measures, it enhances the comprehensiveness of PatentScore, ensuring that both form and meaning are adequately captured.
6 Conclusion
This research proposed PatentScore, a multi-dimensional framework that systematically integrates structural, legal, and semantic metrics for evaluating patent claims. Unlike conventional NLG metrics that fail to capture domain-specific requirements, PatentScore explicitly incorporates legal soundness and structural validity, achieving strong alignment with expert judgments ($r=0.819$) and significantly reducing error rates.
By bridging legal expertise with language modeling, PatentScore provides a reliable, transparent, and practically useful evaluation foundation, establishing a pathway for more trustworthy and innovative applications of AI in the patent domain. The experimental results show that structural and legal factors play a decisive role in assessing claim quality, while semantic metrics provide complementary value by reinforcing terminology consistency. Through expert-based validation, comparative benchmarking, and ablation studies, PatentScore is demonstrated to be a rigorous, interpretable, and reliable tool for evaluating LLM-generated claims.
While the proposed framework is developed with a focus on patent claims, it can be easily extended to other high-precision document types that require structural constraints and legal clarity, such as contract clauses, policy reports, and regulatory compliance in clinical documentation. This contributes to the broader scope of establishing domain-specific evaluation criteria for the trustworthy deployment of LLMs in high-stakes settings.
Limitations
This study has several limitations, which also suggest promising directions for future research.
First, the dataset used in this study is restricted to a specific technical domain and filing period. Expanding the scope to include a broader range of technologies and time frames would enhance the generalizability of the framework.
Second, PatentScore primarily evaluates intrinsic aspects of claim quality, such as structural, legal, and semantic dimensions, but does not address extrinsic properties like novelty or distinctiveness. Future work could incorporate comparative modules that measure overlap with prior art, enabling assessments of uniqueness and scope validity, which are critical in real-world patent examination.
Third, while PatentScore is designed for evaluating LLM-generated claims, its modular architecture allows adaptation to human-written patents as well. This extension would enable automated analysis and feedback on structural coherence, clarity, and legal soundness in expert-authored claims.
Fourth, integrating legal knowledge graphs into the generation and evaluation pipeline could further improve factual accuracy and contextual alignment. Such integration would strengthen the framework’s ability to handle complex dependencies across legal and technical contexts.
Ethical Considerations
In conducting this research, we ensured the highest ethical standards were upheld. All data used in this study, including GPT-4o-mini generated claims and the HUPD dataset, were obtained from publicly available sources and used in compliance with the respective licenses, such as the CC BY License. Human evaluators who participated in the evaluation process provided voluntary and informed consent before contributing to the study. Their identities remain anonymous to protect their privacy and confidentiality.
The study does not involve any personally identifiable information or sensitive data, and it strictly adheres to ethical guidelines for research involving human evaluators. Furthermore, we acknowledge the potential for misuse of AI in automating patent generation and emphasize the importance of developing AI tools that align with legal and ethical standards. The proposed framework, PatentScore, aims to improve transparency, reliability, and accountability in AI-generated patent claims.
Acknowledgments
This work was supported by the Australian Research Council (ARC) under the following projects: LP230201022 (Ethical Enterprise Representations for Personalised Sustainable Finance, Australian Research Council Linkage Grants), DP240102050 (Data Complexity and Uncertainty-Resilient Deep Variational Learning, Australian Research Council Discovery Grants), and LE240100131 (Federated Omniverse Facilities for Smart Digital Futures, Australian Research Council Linkage Infrastructure, Equipment and Facilities Grants).
We thank the creators of the HUPD dataset (CC BY License) and OpenAI for access to GPT, both of which were essential resources for this study. We are also grateful to the three evaluators for their valuable contributions, careful assessments, and voluntary participation after providing informed consent. Their expertise and feedback were instrumental in validating the evaluation framework. Finally, we acknowledge the broader research community for developing the open-source tools and prior works that served as the foundation for this study.
Appendix A Appendix
A.1 Prompts for Experiments
A.1.1 Prompts for Claim Structure Evaluation
Name: Patent Claim Structure
I. Overview
Definition: A patent claim’s composition consisting of three essential parts (preamble, transitional phrase, and elements), which verify whether all these elements are properly positioned.
Evaluation Purpose: To systematically assess the structural completeness and proper organization of a patent claim to ensure clear and enforceable rights.
II. Key Evaluation Points
  1. 1.
    Preamble identification and content validation
  2. 2.
    Transitional phrase appropriateness and placement
III. Important Notes
  • Ensure all three essential parts exist.
  • Check proper element positioning.
IV. Standard Examples
"A device for processing data, comprising: a memory configured to store data; a processor coupled to the memory; and an interface connected to the processor."
V. Evaluation Procedure
  1. 1.
    Content Identification
    Input: Claim 1 requiring evaluation
    Output: Checklist, mapping, and element classification
    Q: Have all structural elements been identified?
    Tasks:
    • Carefully read the provided Claim 1 to identify:
      • Preamble
      • Transitional phrase
      • Elements/components
    • Verify presence of all essential parts:
      • Preamble completeness
      • Appropriate transitional phrase
      • All necessary elements
    • Create mapping of identified elements.
  2. 2.
    Critical Element Verification
    Input: Identified elements and relationships
    Output: Verification report of required elements
    Q: Do the essential elements meet requirements?
    Tasks:
    • Confirm essential structural elements:
      • Clear and complete preamble statement
      • Standard transitional phrase usage
      • Properly structured body elements
    • Verify element relationships:
      • Proper hierarchical structure
    • Remember missing or inappropriate items
  3. 3.
    Format Compliance Assessment
    Input: Element verification results
    Output: Compliance evaluation report
    Q: Does it comply with format requirements?
    Tasks:
    • Check format compliance:
      • Indentation and spacing
      • Punctuation usage (colons, semicolons)
      • Conjunction placement
    • Identify rule violations
    • Remember formatting errors
  4. 4.
    Standard-Based Comparative Analysis
    Input: Compliance assessment results
    Output: Comparative analysis report
    Q: How well does it match standards/examples?
    Tasks:
    • Compare with standard claim structures:
      • Similar technology field examples
      • Accepted structural patterns
    • Identify structural deviations
    • Remember errors
  5. 5.
    Final Scoring
    Input: Analysis results from previous steps
    Output: Final numerical score
    Patent Claim Structure Scoring (1-5 points):
    • 1 point (<20%):
      Example: "Processing data with memory and processor."
      Issues: No clear preamble, no transitional phrase, unstructured elements, missing punctuation, no formatting.
    • 2 points (20-40%):
      Example: "A data processing device having memory for storing data processor for processing interface for connecting."
      Issues: Weak preamble, missing transitional phrase, poor organization, incorrect punctuation, basic elements only.
    • 3 points (41-60%):
      Example: "A device for processing data comprising: a memory that stores data a processor connected to memory an interface; wherein the processor processes data."
      Issues: Basic preamble, has transitional phrase, basic formatting, some punctuation errors, basic relationships.
    • 4 points (61-80%):
      Example: "A device for processing data, comprising: a memory configured to store data; a processor coupled to the memory; an interface connected to the processor; wherein the interface transmits data."
      Strengths: Clear preamble, correct transitional phrase, good formatting, minor structural issues, clear relationships.
    • 5 points (>80%):
      Example: "A device for processing data, comprising: a memory configured to store input data; a processor coupled to the memory and configured to process data; an interface coupled to the processor; wherein the processor updates rules based on feedback."
      Strengths: Perfect structure, proper formatting, clear hierarchy, complete elements, professional composition.
A.1.2 Prompts for Punctuation in Patent Claims
Name: Punctuation in Patent Claim
I. Overview
Definition: Evaluation of proper use and placement of essential punctuation marks in patent claims (colon, semicolons, "and", period).
Evaluation Purpose: To systematically assess the accuracy of punctuation usage to ensure clarity and readability of the claim.
II. Key Evaluation Points
  1. 1.
    Essential punctuation mark validation
  2. 2.
    Appropriate punctuation placement
III. Important Notes
  • Check for colon (:) after transitional phrase.
  • Check for semicolons (;) after elements.
  • Check for "and" before the last element.
  • Check for period (.) at the end of claim.
IV. Standard Examples
"A device for processing data, comprising: a memory configured to store data; a processor coupled to the memory; and an interface connected to the processor."
V. Evaluation Procedure
  1. 1.
    Content Identification
    Input: Claim 1 requiring evaluation
    Output: Checklist, mapping, and punctuation classification
    Q: Have all punctuation marks been identified?
    Tasks:
    • Carefully read the provided Claim 1 to identify:
      • Colon location
      • Semicolon locations
      • "And" location
      • Period location
    • Verify presence of all essential punctuation:
      • Colon after transition
      • Semicolons after each element
      • "And" before last element
      • Period at claim end
    • Create mapping of identified punctuation
  2. 2.
    Critical Element Verification
    Input: Identified punctuation and locations
    Output: Essential punctuation verification report
    Q: Do the essential punctuation marks meet requirements?
    Tasks:
    • Confirm essential punctuation:
      • Appropriate colon usage
      • Appropriate semicolon usage
      • Appropriate "and" usage
    • Verify punctuation locations:
      • Exact position of each mark
    • Remember missing or inappropriate items
  3. 3.
    Format Compliance Assessment
    Input: Punctuation verification results
    Output: Compliance evaluation report
    Q: Does punctuation comply with format requirements?
    Tasks:
    • Check punctuation format compliance:
      • Spacing around punctuation
      • Consistency in usage
    • Remember formatting errors
  4. 4.
    Standard-Based Comparative Analysis
    Input: Compliance assessment results
    Output: Comparative analysis report
    Q: How well does it match standards/examples?
    Tasks:
    • Compare with standard punctuation usage:
      • Standard case patterns
      • Accepted usage practices
    • Identify punctuation deviations
    • Remember errors
  5. 5.
    Final Scoring
    Input: Analysis results from previous steps
    Output: Final numerical score
    Patent Claim Punctuation Scoring (1-5 points):
    • 1 point (<20%):
      Example: "A device with memory processor and interface."
      Issues: No colon after transition, no semicolons between elements, missing "and", missing period, no spacing after marks.
    • 2 points (20-40%):
      Example: "A device comprising memory; processor; interface."
      Issues: Missing colon after transition, inconsistent semicolon use, missing "and", has period, poor spacing.
    • 3 points (41-60%):
      Example: "A device comprising: memory; processor; and interface."
      Issues: Has colon, some semicolons, has "and", missing period, basic spacing.
    • 4 points (61-80%):
      Example: "A device, comprising: memory; processor; and interface."
      Strengths: Proper colon, most semicolons correct, proper "and", has period, minor spacing issues.
    • 5 points (>80%):
      Example: "A device for processing data, comprising: a memory configured to store data; a processor coupled to the memory; and an interface connected to the processor."
      Strengths: Perfect colon placement, all semicolons correct, proper "and" placement, correct period, perfect spacing.
A.1.3 Claim Inconsistent Element Referencing
Name: Claim Inconsistent Element Referencing
I. Overview
Definition: Assessment of consistency in referring to claim components (e.g., touchscreen display, processor, battery), ensuring each element maintains uniform terminology throughout the claim and avoids ambiguous references.
Evaluation Purpose: To ensure each component in the claim is consistently referenced using precise terminology, maintaining clarity and avoiding confusion in element relationships.
II. Key Evaluation Points
  1. 1.
    Terminology consistency verification
  2. 2.
    Reference clarity assessment
III. Important Notes
  • Check for consistent use of component terms
  • Verify clear reference relationships
  • Confirm no ambiguous references
IV. Standard Examples
"A device, comprising: a touchscreen display; a processor connected to the touchscreen display; and a battery powering the processor and the touchscreen display."
V. Evaluation Procedure
  1. 1.
    Content Identification
    Input: Claim 1 requiring evaluation
    Output: Checklist, mapping, and reference classification
    Q: Have all element references been identified?
    Tasks:
    • Carefully read the provided Claim 1 to identify:
      • Each component term
      • Each reference to components
      • Reference relationships
    • Verify component terminology:
      • Initial component definitions
      • Subsequent references
      • Reference consistency
    • Create mapping of identified references
  2. 2.
    Critical Element Verification
    Input: Identified references and relationships
    Output: Reference verification report
    Q: Do the reference terms meet requirements?
    Tasks:
    • Confirm reference consistency:
      • Uniform terminology use
      • Clear reference connections
      • Proper reference terms
    • Verify reference accuracy:
      • Correct component references
    • Remember inconsistent or improper items
  3. 3.
    Format Compliance Assessment
    Input: Reference verification results
    Output: Compliance evaluation report
    Q: Does reference usage comply with format requirements?
    Tasks:
    • Check reference format compliance:
      • Reference term consistency
      • Reference terminology clarity
    • Remember formatting errors
  4. 4.
    Standard-Based Comparative Analysis
    Input: Compliance assessment results
    Output: Comparative analysis report
    Q: How well does it match standards/examples?
    Tasks:
    • Compare with standard reference patterns:
      • Standard reference cases
      • Accepted reference practices
    • Identify reference deviations
    • Remember errors
  5. 5.
    Final Scoring
    Input: Analysis results from previous steps
    Output: Final numerical score
    Patent Claim Element Reference Consistency Scoring (1-5 points):
    • 1 point (<20%):
      Example: "A device including a screen connected to CPU, wherein the display is connected to processor, and a battery powering the touchscreen."  Issues: Different terms for same element (screen/display/touchscreen), incorrect reference terms (CPU/processor), unclear reference relationships, serious inconsistencies, confusing element relationships.
    • 2 points (20-40%):
      Example: "A touchscreen display; a processor connected to the screen; and a battery powering the CPU and display."
      Issues: Some inconsistent terminology, mixed reference terms, unclear reference relationships, major inconsistencies, basic element mentions only.
    • 3 points (41-60%):
      Example: "A touchscreen display; a processor connected to the touchscreen display; and a battery powering the processor and display."  Issues: Basic consistency, some missing references, partially clear relationships, some inconsistencies, basic reference structure.
    • 4 points (61-80%):
      Example: "A touchscreen display; a processor connected to said touchscreen display; and a battery powering said processor and said touchscreen display."
      Strengths: Mostly consistent terminology, proper reference terms, clear reference relationships, minor inconsistencies, clear element relationships.
    • 5 points (>80%):
      Example: "A device, comprising: a touchscreen display; a processor operatively connected to said touchscreen display; and a battery configured to power said processor and said touchscreen display."
      Strengths: Perfect terminology consistency, accurate reference terms, clear reference relationships, professional composition, perfect element relationships.
A.1.4 Patent Claim Antecedent Basis
Name: Patent Claim Antecedent Basis
I. Overview
Definition: Assessment of proper introduction and subsequent referencing of claim elements, ensuring each "the" reference has a corresponding previous introduction with "a" or "an".
Evaluation Purpose: To ensure clear and legally proper element referencing by confirming proper establishment of antecedent basis for all claim elements.
II. Key Evaluation Points
  1. 1.
    Article usage verification (a, an, the)
  2. 2.
    First mention validation
III. Important Notes
  • Check first introduction uses "a" or "an"
  • Verify subsequent references use "the"
  • Confirm all "the" references have prior introduction
IV. Standard Examples
"A device, comprising: a display; a processor connected to the display; and a battery, wherein the battery powers the processor and the display."
V. Evaluation Procedure
  1. 1.
    Content Identification
    Input: Claim 1 requiring evaluation
    Output: Checklist, mapping, and article usage classification
    Q: Have all article usages been identified?
    Tasks:
    • Carefully read the provided Claim 1 to identify:
      • First mentions (a, an)
      • Subsequent references (the)
      • Reference relationships
    • Verify article usage:
      • Initial introductions
      • Subsequent references
      • Reference consistency
    • Create mapping of identified articles.
  2. 2.
    Critical Element Verification
    Input: Identified articles and relationships
    Output: Article usage verification report
    Q: Do the article usages meet requirements?
    Tasks:
    • Confirm proper introduction:
      • Proper first mention articles
      • Proper subsequent articles
    • Verify reference accuracy:
      • Match between references
    • Remember improper article usage
  3. 3.
    Format Compliance Assessment
    Input: Article verification results
    Output: Compliance evaluation report
    Q: Does article usage comply with requirements?
    Tasks:
    • Check article usage compliance:
      • Article sequence
      • Reference clarity
    • Remember formatting errors
  4. 4.
    Standard-Based Comparative Analysis
    Input: Compliance assessment results
    Output: Comparative analysis report
    Q: How well does it match standards/examples?
    Tasks:
    • Compare with standard article usage:
      • Standard reference patterns
      • Accepted article practices
    • Identify article usage deviations
    • Remember errors
  5. 5.
    Final Scoring
    Input: Analysis results from previous steps
    Output: Final numerical score
    Patent Claim Antecedent Basis Scoring (1-5 points):
    • 1 point (<20%):
      Example: "The display and the CPU, wherein the memory connects to the processor."
      Issues: Improper first mentions, references without antecedent basis, incorrect article usage, severe antecedent basis issues, confusing relationships.
    • 2 points (20-40%):
      Example: "The display and a processor, wherein the CPU connects to memory."
      Issues: Some improper first mentions, incomplete antecedent basis, incorrect article usage, major antecedent basis issues, basic elements only.
    • 3 points (41-60%):
      Example: "A display; the processor connected to the display; and the memory connected to the processor."
      Issues: Basic antecedent basis, some incorrect articles, some unclear references, some antecedent basis issues, basic reference structure.
    • 4 points (61-80%):
      Example: "A display; a processor connected to the display; and the processor connected to a memory."
      Strengths: Mostly proper antecedent basis, proper article usage, clear references, minor antecedent basis issues, clear element relationships.
    • 5 points (>80%):
      Example: "A display; a processor connected to the display; a memory; and a battery connected to the processor and the memory."
      Strengths: Perfect antecedent basis, accurate article usage, clear references, professional composition, perfect element relationships.
A.1.5 Claim Validity & Uniqueness
Name: Claim Validity & Uniqueness
I. Overview
Definition: Evaluation of whether each claim component contributes new technical features and specific elements to the invention, without contradictions or unnecessary repetitions of earlier components.
Evaluation Purpose: To ensure the claim maintains its validity and uniqueness by differentiating each part and upholding legal strength within the patent framework.
II. Key Evaluation Points
  1. 1.
    Technical feature uniqueness
  2. 2.
    Component differentiation
III. Important Notes
  • Verify each component adds unique value
  • Check for technical contradictions
  • Confirm no unnecessary repetitions
IV. Standard Examples
"A data processing device, comprising: a unique processing unit having specific computational capabilities; a specialized memory unit configured to enhance processing speed; and a novel interface designed to optimize data transfer."
V. Evaluation Procedure
  1. 1.
    Content Identification
    Input: Claim 1 requiring evaluation
    Output: Checklist, mapping, and uniqueness classification
    Q: Have all unique elements been identified?
    Tasks:
    • Carefully read the provided Claim 1 to identify:
      • Technical features
      • Unique elements
      • Specific contributions
    • Verify element uniqueness:
      • Technical differentiation
      • Specific features
    • Create mapping of unique features
  2. 2.
    Critical Element Verification
    Input: Identified elements and features
    Output: Validity verification report
    Q: Do the elements contribute unique value?
    Tasks:
    • Confirm technical contribution:
      • Novel features
      • Specific improvements
    • Verify element differentiation:
      • Technical distinctions
    • Remember contradictions or repetitions
  3. 3.
    Format Compliance Assessment
    Input: Verification results
    Output: Compliance evaluation report
    Q: Does the claim maintain validity requirements?
    Tasks:
    • Check validity compliance:
      • Technical consistency
      • Feature uniqueness
    • Remember validity issues
  4. 4.
    Standard-Based Comparative Analysis
    Input: Compliance assessment results
    Output: Comparative analysis report
    Q: How well does it establish uniqueness?
    Tasks:
    • Compare with standard features:
      • Similar technology features
      • Distinctive elements
    • Identify uniqueness deviations
    • Remember errors.
  5. 5.
    Final Scoring
    Input: Analysis results from previous steps
    Output: Final numerical score
    Patent Claim Validity & Uniqueness Scoring (1-5 points):
    • 1 point (<20%):
      Example: "A data processing device comprising a processor and memory."
      Issues: No technical features, generic components only, lack of differentiation, no uniqueness, vague implementation.
    • 2 points (20-40%):
      Example: "A data processing device with a high-speed processor and large memory."
      Issues: Minimal technical features, basic differentiation attempt, insufficient specificity, limited uniqueness, common technical level.
    • 3 points (41-60%):
      Example: "A data processing device comprising: a parallel processing processor; a high-speed cache memory; and a data optimization interface."
      Issues: Basic technical features, some differentiation, basic uniqueness, partial specificity, some technical improvements.
    • 4 points (61-80%):
      Example: "A data processing device comprising: a multi-core parallel processor; a hierarchical cache memory structure; and a real-time data optimization interface."
      Strengths: Clear technical features, specific differentiation, good uniqueness, concrete implementation, clear technical improvements.
    • 5 points (>80%):
      Example: "A data processing device comprising: an AI-accelerated multi-core processor; a dynamic memory allocation hierarchical cache; and an intelligent interface with adaptive data optimization algorithms."
      Strengths: Strong technical features, excellent differentiation, high uniqueness, detailed implementation, innovative improvements.
A.1.6 Ambiguous Claim Scope
Name: Ambiguous Claim Scope
I. Overview
Definition: Evaluation of claim language clarity and boundary definition, ensuring specific and unambiguous protection scope without unclear or overly broad terms that could lead to uncertainty in claim interpretation.
Evaluation Purpose: To ensure the claim clearly defines its protection boundaries using specific and precise language, avoiding ambiguous or overly broad terms that could weaken legal protection.
II. Key Evaluation Points
  1. 1.
    Boundary clarity verification
  2. 2.
    Term specificity assessment
III. Important Notes
  • Check for clear scope boundaries
  • Verify specific technical terms
  • Confirm no ambiguous language
IV. Standard Examples
"A data processing device, comprising: a 2.4 GHz processor specifically configured for image processing; a 512GB solid-state memory connected to the processor; and a 4K resolution display interface operating at 60Hz refresh rate."
V. Evaluation Procedure
  1. 1.
    Content Identification
    Input: Claim 1 requiring evaluation
    Output: Checklist, mapping, and scope classification
    Q: Have all scope-defining elements been identified?
    Tasks:
    • Carefully read the provided Claim 1 to identify:
      • Scope boundaries
      • Technical specifications
      • Limiting terms
    • Verify term specificity:
      • Technical parameters
      • Functional limitations
    • Create mapping of scope elements
  2. 2.
    Critical Element Verification
    Input: Identified scope elements and boundaries
    Output: Scope verification report
    Q: Do the elements clearly define boundaries?
    Tasks:
    • Confirm boundary clarity:
      • Clear technical limits
      • Specific definitions
    • Verify scope precision:
      • Term specificity
    • Remember ambiguous or broad terms
  3. 3.
    Format Compliance Assessment
    Input: Scope verification results
    Output: Compliance evaluation report
    Q: Does the scope comply with clarity requirements?
    Tasks:
    • Check scope compliance:
      • Term precision
      • Boundary clarity
    • Remember clarity issues
  4. 4.
    Standard-Based Comparative Analysis
    Input: Compliance assessment results
    Output: Comparative analysis report
    Q: How well does it define protection scope?
    Tasks:
    • Compare with standard scope definitions:
      • Similar technology scopes
      • Accepted boundary definitions
    • Identify scope deviations
    • Remember errors
  5. 5.
    Final Scoring
    Input: Analysis results from previous steps
    Output: Final numerical score
    Patent Claim Scope Ambiguity Scoring (1-5 points):
    • 1 point (<20%):
      Example: "An improved device that processes data better."
      Issues: Highly ambiguous terms, unclear boundaries, unmeasurable improvements, subjective expressions, boundless protection scope.
    • 2 points (20-40%):
      Example: "An enhanced processing device with a high-speed processor and large memory."
      Issues: Vague technical terms, broad scope, unclear parameters, generic expressions, wide protection scope.
    • 3 points (41-60%):
      Example: "A processing device comprising: a processor operating above 1GHz; and memory exceeding 256GB."
      Issues: Basic clarity, some specific parameters, basic boundary definition, measurable features, limited protection scope.
    • 4 points (61-80%):
      Example: "A processing device comprising: a 2.4GHz multi-core processor; a 512GB SSD; and a 4K resolution display."
      Strengths: Clear technical terms, specific parameters, clear boundaries, precise measurements, defined protection scope.
    • 5 points (>80%):
      Example: "A processing device comprising: a 2.4GHz 8-core processor performing 1M operations/second; a 512GB NVMe SSD with 3000MB/s read speed; and a 4K 60Hz HDR display."
      Strengths: Very precise technical terms, clear performance parameters, exact boundary definitions, specific measurement criteria, precise protection scope.
Appendix B Prompts for generating by LLM
You are an experienced patent attorney with a deep understanding of patent law and intellectual property (IP) protection. I will provide you with the title, abstract, and detailed description of a specific invention. Based on this information, please draft Claim 1 of the patent, ensuring that it accurately captures the essence of the invention while being legally strong and precise. Claim 1 should encompass the core technology of the invention while meeting legal requirements to distinguish it from prior art. It should use a clear and specific terminology to define the scope of protection appropriately, avoiding overly narrow or unnecessarily broad expressions.
Additionally, the structure of Claim 1 should adhere to the conventional format used in patent legal documents and incorporate terminology commonly used in the relevant technical field.
The Claim 1 you draft must meet the following criteria:
1. Clarity: Avoid ambiguous expressions and use precise language that can be easily understood. 2. Legal Validity: Comply with patent examination standards and be legally appropriate to provide substantial protection. 3. Breadth & Differentiation: Adequately cover the core elements of the invention while highlighting its distinction from existing technology. 4. Technical Accuracy: Accurately reflect the provided invention description without technical errors. 5. Structural Integrity: Follow proper grammar and format for patent claims and maintain logical consistency.
Now, I will provide the patent title, abstract, and detailed description. Based on this information, please draft Claim 1 of the patent.